### v2

# Solving the Inverse Problem of Electrocardiography for Cardiac Digital Twins: A Survey

Abstract
Cardiac digital twins are personalized virtual representations used to understand complex heart mechanisms. Solving the ECG inverse problem is crucial for accurate virtual heart modelling, enabling the derivation of internal electrical activity information from recorded surface potentials. Despite challenges from cardiac complexity, noisy ECG data, and computational efficiency, recent advancements hold significant promise for enhancing virtual heart modelling, ultimately advancing precision medicine in cardiology. This paper aims to provide a comprehensive review of the methods of solving ECG inverse problem, the validation strategies, the clinical applications, and future perspectives. For the computing methodologies, we broadly classify state-of-the-art approaches into two categories: deterministic and probabilistic methods, including conventional and deep learning-based techniques. Integrating physics laws with deep learning models holds promise, but challenges such as capturing dynamic electrophysiology accurately, accessing accurate domain knowledge, and quantifying prediction uncertainty persist. Integrating models into clinical workflows while ensuring interpretability and usability for healthcare professionals is essential. Overcoming these challenges will drive further research in cardiac digital twins.
{IEEEkeywords}
Cardiac electrophysiology, inverse problem, computational model, cardiac digital twin, survey.
1 Introduction
Electrocardiogram (ECG) is a non-invasive technique used to measure the electrical activity of the heart and identify potential abnormalities [1]. It consists of a number of electrodes placed on the skin of the patient, typically on the chest, arms, and legs. These electrodes record the changes in electrical potential on the body surface, which are generated by the underlying electrical activity of the heart. Virtual modelling of cardiac electrophysiology (EP) for digital twins has shown great promise in numerous clinical applications, such as the stratification of arrhythmia risk [2] and the ablation guidance of persistent atrial fibrillation (AF) [3]. However, obtaining numerical solutions to the governing partial differential equations (PDEs) in cardiac EP (ECG forward problem) is quite challenging. Several forward physical modelling approaches have been employed, such as the bidomain model, the monodomain model, and the Eikonal equation. The biodomain model provides information about the ionic currents [4], the monodomain model describes the electrical propagation [5], and the Eikonal equation models wavefront propagation [6]. To obtain a complete understanding of the electrical activity of the heart, we need to estimate the activity inside the heart from the measured ECG signals, namely ECG inverse problem [7]. Fig. 1 presents the inverse inference procedure for reconstructing the cardiac activity from ECG data. By solving the inverse problem, one can obtain valuable information about the function of the heart, which can aid in the diagnosis and monitoring of various cardiac conditions, as well as the development of personalized treatments for cardiac disorders.
However, directly predicting cardiac activity from standard ECG signals (often recorded using a limited number of electrodes, such as the 12-lead ECG with ten electrodes) is a challenging task. This is mainly because the limited number of electrodes in standard ECGs makes it difficult to capture the fine-grained details of cardiac electrodynamics across the entire heart. Therefore, body surface potential (BSP) mapping has been developed by deploying a larger number of ECG electrodes, such as Dalhousie Standard Torson 120 electrodes, to provide high-resolution spatiotemporal information about cardiac electrodynamics projected onto the body surface [8]. It allows for detailed mapping and localization of cardiac activity and thus has been widely used in the last decades for solving the ECG inverse problem. Nevertheless, the ECG inverse problem, like many other inverse problems, faces two primary challenges. Firstly, the problem is ill-posed, meaning that it may not have a unique solution and small errors in the input data can lead to significant errors in the output. Secondly, solving the ECG inverse problem can be computationally intensive, particularly when dealing with high-resolution BSP mapping data. The complex geometrical, mechanical, and electrical properties of the heart, as well as the need for real-time or near-real-time solutions in clinical applications, pose computational challenges.
Figure 1: Illustration of the electrocardiogram (ECG) inverse problem.
Table 1: Search engines and expressions used to identify potential papers for review.
Engine Google Scholar, PubMed, IEEE-Xplore, and Citeseer
Term “Electrocardiogram” or “ECG” or “EKG” or “electrocardiographic” or “electrophysiological” or “electrophysiology” or “EP” or “electrocardiography” or “body surface potential mapping” or “BSPM” and
“Inverse problem/ issue” or “inverse solution/ estimation/ inference/ reconstruction” or “cardiac model personalization” or “electrocardiographic imaging/ mapping” or “ECG-imaging” or “ECGI” or “activation imaging” or “noninvasive imaging” and
“Electrical activity/ excitation” or “surface potential” or “epicardial/ endocardial potential” or “action potential” or “transmembrane potential” or “activation sequence” or “activation wavefront” or “activation time map” or “pacing site” or “atrial potential” or “local parameter”
Figure 2: Timeline of the development of ECG, imaging, and cardiac digital twins.
1.1 History: From Electrocardiography to Cardiac Digital Twins
Investigations of cardiac electrophysiology can be traced back to 1842 when Carlo Matteucci identified signs of cardiac electrical activity in frog hearts [9]. Dr. Augustus Waller furthered this research by recording the first human electrocardiogram in 1887, using a capillary electrometer and electrodes placed on the chest and back of a human subject [10]. His study demonstrated that electrical activity preceded ventricular contraction. Subsequently, in 1893, Einthoven coined the term “electrocardiogram” to describe the cardiac waveforms [11]. The American Heart Association (AHA) standardized the 12-lead ECG in 1954 [12], paving the way for advancements in cardiac electrophysiology. In 1960, Denis Noble developed one of the first mathematical models of a cardiac electric cell based on Hodgkin-Huxley equations, further refining it to characterize the long-lasting action and pacemaker potentials observed in the Purkinje fibres of the heart [13, 14, 15]. This laid the groundwork for cardiac simulations using computer models. Meanwhile, body surface potential mapping (BSPM) emerged as an alternative method for non-invasive and high-resolution measurement of human cardiac electrical activity in 1963 [16]. In 1970, Durrer et al. [17] conducted a seminal investigation into the electrical behavior of the human heart at the organ level. Since 1972, numerous studies have tackled the electrocardiography inverse problem using analytical and computational models [18, 19, 20, 21], experimental animals [22, 23], and human subjects [24, 25]. Advancements in medical imaging have also enabled non-invasive visualization of cardiac anatomy since the inception of techniques like the first human MRI scan in 1977 [26]. The formalization of inverse procedures reconstructing bioelectric sources from BSPM and medical imaging led to the development of electrocardiographic imaging (ECGi) techniques, notably pioneered by Yoram Rudy’s lab [27, 28, 29, 30, 31].
The concept of the digital twin was formally introduced in 2002 by Michael Grieves, though it can be traced back to NASA’s use for the Apollo 13 rescue mission in 1970 [32]. The Physiome Project was initiated in 2003 with the aim of establishing a framework for modelling the human body, utilizing computational methods to integrate biochemical, biophysical, and anatomical information across cells, tissues, and organs [33]. Starting from 2014, many companies like Dassault, Siemens, and ANSYS embraced the term “digital twins” in their marketing initiatives. For example, Dassault Systèmes initiated the Living Heart Project in 2014, which is the first comprehensive computer model of the human heart, integrating all functional aspects, i.e., blood flow dynamics, mechanical properties, and electrical conduction [34]. In recent years, digital twin tools have become popular in healthcare, such as cardiac digital twins [35], brain digital twins [36], etc. In 2016, the US Food & Drug Administration (FDA) Center for Devices and Radiological Health released the first guidance to enable the use of modelling and simulation to develop in silico trials [37]. In 2019, Trayanova’s lab received FDA approval for a randomized clinical trial involving 160 patients, termed OPTIMA (OPtimal Target Identification via Modeling of Arrhythmogenesis), to demonstrate the effectiveness of cardiac digital twins in guiding atrial ablation procedures [3]. The brief timeline has been presented in Fig. 2.
1.2 Study Inclusion and Literature Search
In this work, we aim to provide readers with a survey of the state-of-the-art solutions to electrocardiographic inverse problems. There exist several modalities for cardiac electrical activity assessment, such as ECG, BSP mapping, electrograms (EGM), electro-anatomic mapping (EAM), vectorcardiography (VCG), etc. However, in the context of solving the ECG inverse problem, we usually only involve ECG and BSP as the observed data. Note that compared to standard 12-lead ECG, BSP mapping is, however, not yet widely available as a diagnostic modality [38]. For cardiac activity representation, two major source models are employed, i.e., heart surface potential (HSP) and transmembrance potential defined on the volumetric mesh of the heart. Specifically, it can be further categorized into epicardial and/ or endocardial potential, activation time map/ activation wavefront, transmembrane potential, action potential duration map, voltage potential maps, isochrones/ isochrone maps, etc.
Table 2: Summary of the related representative existing surveys. DL: deep learning; MCG: Magnetocardiography; ATM: activation time maps; AF: atrial fibrillation.
Source & Publish Year Scope Inclusion
Macleod et al. (1998) [7] ECG inverse problem conventional solutions
Dossel et al. (2000) [39] ECG and MCG inverse problem conventional solutions
Yaman et al. (2013) [40] inverse problem for applied sciences not for ECG; conventional solutions
Uhlmann et al. (2014) [41] inverse problem not for ECG; conventional solutions
Calvetti et al. (2018) [42] inverse problem not for ECG; conventional solutions
Genzel et al. (2022) [43] inverse problem not for ECG; mainly DL-based solutions
Arridge et al. (2019) [44] inverse problem not for ECG; mainly DL-based solutions
Sameni et al. (2010) [45] fetal ECG analysis fetal ECG; partially includes inverse problem
Bear et al. (2015) [46] ECG forward problem not for inverse problem
Hong et al. (2020) [47] DL-based ECG analysis not for inverse problem
Liu et al. (2021) [48] DL-based ECG analysis not for inverse problem
Somani et al. (2021) [1] DL-based ECG analysis not for inverse problem
Bifulco et al. (2021) [49] translational applications of computational modelling translational applications
Cantwell et al. (2015) [50] ECG inverse problem mainly focusing on the inference of ATM and CV
Cluitmans et al. (2015) [51] ECG inverse problem mainly focusing on conventional methods
Cluitmans et al. (2018) [52] ECG inverse problem mainly focusing on the validation and clinical application
Yadan et al. (2023) [53] ECG inverse problem mainly focusing on the identification of AF drivers
Hernandez et al. (2023) [54] ECG inverse problem mainly focusing on the atria
To ensure comprehensive coverage, we have screened publications mainly from the last 24 years (2000-2024) related to this topic. Our main sources of reference were Internet searches using engines, including Google Scholar, PubMed, IEEE-Xplore, Connected Papers, and Citeseer. To cover as many related studies as possible, flexible search terms have been employed when using these search engines, as summarized in Table 1. There are several research teams and industry companies at the forefront of computational cardiology and electrocardiography, working to advance our understanding of cardiac function, arrhythmias, and heart-related disorders through computational methods and simulations  [^1] Continuously updating compilation of cardiac digital twin resources: github.com/lileitech/Awesome-Cardiac-Digital-Twins. We followed their recent works to obtain state-of-the-art methodologies and clinical applications. Both peer-reviewed journal papers and conference papers were included here. In the way described above, we have collected a comprehensive library of more than 200 papers. Note that we picked the most detailed and representative ones for this review when we encountered several papers from the same authors about the same subject.
1.3 Existing Survey from Literature
Table 2 lists current review papers related to the topic, i.e., ECG inverse problem. One can see that the scopes of current review works are different from ours, though with partial overlaps. For example, there are several existing reviews focusing on summarizing solutions of inverse problems but not for ECG data [40, 41, 42, 44, 43]. Alternatively, there exist few surveys on the ECG analysis, but not focusing on the inverse problem [46, 48, 1]. Bifulco et al. [49] provided the potential clinical applications of the ECG inverse inference strategies, but it is not a methodological survey. Although Macleod et al. [7] and Dossel et al. [39] performed a methodology survey on the ECG inverse problem, we focus on the developments proceeding their reviews. Cantwell et al. [50] mainly summarized the techniques for local activation time annotation and conduction velocity estimation in cardiac mapping. We provide a comprehensive and up to date review of ECG inverse inference, encompassing both its recent methodological aspects and potential clinical applications.
Figure 3: Summary of the ECG inverse inference methodologies. MCMC: Markov Chain Monte Carlo.
2 Methodology
2.1 Definition of the Problem
The ECG inverse problem refers to the mathematical task of determining the electrical activity within the heart based on measurements of the ECG signals recorded on the body surface. It involves applying a quasi-static approximation of Maxwell’s equations on the electromagnetic field between the heart and the torso to establish a relationship between the ECG data and the cardiac electrical excitation. Solving these equations numerically on subject-specific heart-torso models yields a linear measurement model,
$\boldsymbol{y}(\boldsymbol{s},t)=\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)+\epsilon,$ (1)
where $\boldsymbol{y}(\boldsymbol{s},t)$ and $\boldsymbol{u}(\boldsymbol{s},t)$ refer to the vector of ECG signals and cardiac electrical parameters at position $\boldsymbol{s}$ and time instant $t$, $\boldsymbol{H}$ denotes the forward matrix, and $\epsilon$ represents the noise term that accounts for measurement and modelling errors. Based on this physical system modelled by PDEs and/or ordinary differential equations (ODEs), the typical solution of the problem is to inversely estimate $\boldsymbol{u}(\boldsymbol{s},t)$ that minimizes the fitting of the observational data $\boldsymbol{y}(\boldsymbol{s},t)$,
$\min_{\boldsymbol{u}(\boldsymbol{s},t)}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2},$ (2)
which however is ill-posed. The inverse problem can exhibit one of three characteristics: it may have multiple solutions (non-unique), no solution (non-existence), or solutions that are highly sensitive to small perturbations in the measured data (unstable) [55]. Moreover, it is typically influenced by measured input data that is frequently corrupted by random or stochastic noise.
To solve the ECG inverse problem, the state-of-the-art approaches can be coarsely separated into two kinds: deterministic and probabilistic methods [56], as presented in Fig. 3. Deterministic approaches in cardiac electrophysiology involve minimizing a cost function that quantifies the discrepancy between the observed data and the model predictions. Rather than aiming solely to minimize a cost function, probabilistic models embrace uncertainty and provide a probabilistic distribution over possible outcomes.
2.2 Deterministic Estimation Approaches
2.2.1 Spatial Regularization
To stabilize and improve the solution of the ECG inverse problem, spatial regularization methods are employed [57]. These methods introduce spatial constraints to promote smoothness in the estimated electrical activities across the heart surface at different orders of derivatives [58]. The commonly used spatial regularization methods include Tikhonov regularization [24], truncated singular value decomposition (TSVD) [59], truncated total least squares (TTLS) [60], L1-norm regularization [61], and total variation (TV) regularization [62]. These approaches aim to minimize the impact of noise and improve the accuracy of the inverse solution. For example, the Tikhonov regularization, including both zero-order and first-order variants, is the most widely used regularization technique [24, 63, 64, 65, 66, 67]. It minimizes the mean squared error from the body-heart transformation while simultaneously penalizing the L2 norm of the inverse HSP solution. The objective function can be formulated as,
$\min_{\boldsymbol{u}(\boldsymbol{s},t)}\left\{\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}+\lambda_{Tikh}^{2}\|\Gamma\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}\right\},$ (3)
where the notation $\|\cdot\|^{2}$ represents the L2 norm, $\lambda_{Tikh}$ denotes the regularization coefficient, and $\Gamma$ is the operator constraining the HSP solution $u(\mathbf{s},t)$, which is an identity matrix for zero-order regularization and a spatial gradient operator for first-order regularization to increase the spatial smoothness of the inverse solution. This dual process effectively suppresses the unreliable components in the solution and enhances overall smoothness for improved accuracy. TSVD regularization method directly filters small singular values of the transfer matrix $\boldsymbol{H}$ and thus be robust to noises. TTLS regularization is effective in processing geometric errors that exist in the ECG inverse problem [60]. L1-norm and TV regularization techniques are useful for addressing specific challenges in the inverse problem, such as noise suppression, sparsity, and edge preservation.
2.2.2 Spatiotemporal Regularization
While spatial regularization methods are effective and widely used in ECG inverse problems, they often treat the problem as a static spatial problem and ignore the temporal dynamics of cardiac electrical activity. In some cases, this assumption may be reasonable, especially for certain types of cardiac abnormalities or when the temporal variations in the electrical activity are relatively slow compared to the sampling rate. However, in situations where the temporal dynamics are critical or when the electrical activity changes rapidly, temporal regularization methods that explicitly consider the time-varying nature of the problem may be more appropriate. These methods can help capture dynamic changes in electrical activity over time and provide more accurate and physiologically relevant solutions.
To fuse the space-time correlations into the inverse ECG problem, spatiotemporal regularizations are employed for robust estimation of HSP. For instance, besides employing Tikhonov regularization in the spatial domain, Messnarz et al. [68] introduced additional temporal constraints by assuming that cardiac electric potentials are non-decreasing during the depolarization phase. Yao et al. [69] proposed a spatiotemporal regularization method by regularizing both the spatial and temporal smoothness in the inverse ECG modelling, with the following objective function,
$\min_{\boldsymbol{u}(s,t)}\sum_{t}\bigg{\{}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}+\lambda_{s}^{2}\|\Gamma\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}$ (4)
$+\lambda_{t}^{2}\sum_{\tau=t-\omega/2}^{\tau=t+\omega/2}\|\boldsymbol{u}(\boldsymbol{s},t)-\boldsymbol{u}(\boldsymbol{s},\tau)\|_{2}^{2}\bigg{\}},$
where $\lambda_{s}$ and $\lambda_{t}$ are the spatial and temporal regularization terms, respectively, and $\omega$ denotes the selected time window to incorporate the temporal correlation. Splines-based methods can use spline functions to represent the electrical activity on the heart surface over time. They take advantage of the smoothness properties of splines to regularize the inverse solution, effectively capturing both spatial and temporal variations in the electrical potentials [70]. A Kalman filter reformulation can incorporate temporal information to regularize the inverse problem [71, 72]. Recently, to achieve regularization over space and time simultaneously, Cluitmans et al. [73] firstly applied regularization in the sparse wavelet domain for solving the ECG inverse problem, namely wavelet-promoted spatiotemporal regularization. Their experimental results indicate that the sparse representation of the cardiac source is beneficial for improving reconstruction accuracy in electrocardiographic imaging.
2.2.3 Model-based Regularization
Most existing regularization methods primarily rely on the optimization of the given observations, which are normally corrupted by noise in practice. In contrast, model-based regularization can integrate prior physiological knowledge about the electrical propagation inside the heart to constrain the solution space [74]. For instance, step jump functions and logistic functions have been employed to model the activation of action potential [75, 76]. Parameterized curves have been utilized to represent the wavefront velocity as trigonometric functions, while the potential has been modelled using the step response of a second-order linear system [72]. The ECG inverse problems can also be formulated via level-sets by evolving a boundary from an initial region to minimize a filtered residual error, allowing for iterative refinement of the solution. The major advantage of level-set-based methods is that no isotropy assumptions are required, allowing for more flexible modelling and accurate representation of complex geometries and phenomena within the heart [77]. To estimate transmural potentials throughout the myocardium, 3D biophysical EP simulation models have been employed to offer spatiotemporal constraints of the inverse issue [78, 79, 80].
While model-based regularization is effective in regularizing the ill-posed problem, their employed physiological models are developed for the general population [79, 74]. Specifically, the parameters of these models, which are typically linked to tissue characteristics (excitability and contractility of the heart muscle), are often assigned values commonly used in the literature. The general physiological models could overly constrain subject-specific information, leading to potential model errors and model-data mismatch [81]. For instance, the parameters controlling the configuration of the transmembrane potential (TMP) will differ throughout the heart, depending on whether the underlying tissue is in a healthy or diseased state. Therefore, Erem et al. [82] investigated how the solution changes when the assumption of uniform TMP amplitudes throughout the heart fails in the presence of infarction. Similarly, Xu et al. [81] and Dhamala et al. [83] emphasized the importance of uncertainty estimation in providing interpretable reconstructions of cardiac electrical activity.
2.2.4 Data-Driven Neural Networks
The notable rise in computational resources and the vast increase in available data have facilitated the development of data-driven models. Conventional deterministic models for ECG inverse problems typically rely on physics-based mathematical models, such as the bidomain, the monodomain, or the Eikonal model, to describe the propagation of electrical activity through the heart. In contrast, data-driven deterministic models can employ neural networks to directly learn the mapping from ECG data to the underlying electrical sources without relying on explicit physics-based equations. These models are highly flexible and can capture complex, non-linear relationships in the data. They do not rely on strong assumptions about the underlying physiology, making them suitable for cases where the true underlying model is not well-known, varies across patients, or displays unexpectedly abnormal behavior. Therefore, they have been employed to solve the ECG inverse problem. For example, Meister et al. [84] designed a deep learning method based on graph convolutional neural networks (CNNs) to estimate the depolarization patterns in the myocardium with scars. By leveraging the inherent graph structure of the cardiac anatomy, graph CNNs enable the extraction of spatial dependencies and patterns, facilitating a more nuanced representation of electrocardiographic data. Tenderini et al. [85] utilized a PDE-aware deep learning model for epicardial potential (EpiP) reconstruction from ECG data. Camara et al. [86] converted the BSP into an image represented by a 3D matrix and then employed CNNs to predict the locations of AF drivers from BSP. Neumann et al. [87] employed a reinforcement learning model to achieve multi-physics computational model personalization. Similarly, the time delay neural network can employ both current and previous timestep values of BSP to predict the HSP [88, 89, 90]. However, these purely data-driven models generally lack interpretability and may underperform in scenarios with limited data due to their data-intensive training requirements.
Figure 4: Example of the physic-informed neural networks for the reconstruction of activation time map. Here, the Eikonal equation has been employed as an additional physical constraint [91]. Image modified from Sahli [91] with permission.
Figure 5: Example of the physics-based model based on graph convolutional neural networks (CNNs) for the inverse inference of epicardial potential from body surface potential (BSP) [92]. Illustrations designed referring to Jiang et al. [92] and Bear et al. [93].
2.2.5 Physics-Informed Neural Networks
Physics-informed neural networks (PINNs) are data-efficient as they leverage both available data and the known physics equations [94]. They have recently been used to solve inverse problems with hidden physics, including the ill-posed ECG inverse problems [95]. This can be achieved either via introducing underlying physical constraints as an additional loss [95, 96], or designing specialized network architectures that satisfy the physics laws [92, 97]. For instance, Xie et al. [95] combined the data-driven loss and physics-based loss to inversely predict the heart-surface potentials from BSP and evaluated their framework in a 3D torso-heart geometry. Ye et al. [96] proposed a spatial-temporally adaptive PINN framework for solving both ECG forward and inverse problems over 3D bi-ventricular models. Sahli et al. [98] used PINNs in forward mode to estimate activation time maps (ATMs) and conduction velocity (CV) maps, as shown in Fig. 4. They incorporated a physics-informed regularization loss prescribed by the Eikonal equation, which describes the relationship between ATMs and the spatial gradient of CV. Similarly, Herrero et al. [97] developed a PINN model for action potential reconstruction and EP parameter estimation from sparse amounts of EP data. Although they did not inversely infer ATMs from BSP or ECG, a similar idea can be employed to solve the ECG inverse inference problem. Instead of introducing additional loss, Jiang et al. [92] proposed a spatial-temporal graph CNN-based non-Euclidean encoding-decoding model, where the geometry-dependent physics between BSP and HSP can be explicitly modelled via a bipartite graph over their graphical embeddings, as presented in Fig. 5. In general, the use of PINN in cardiac EP is mainly limited to 1D/ 2D in silico data, as it could be computationally expensive in realistic 2D/ 3D settings [97, 96].
2.3 Probabilistic Estimation Approaches
Unlike deterministic regularization techniques, probabilistic or stochastic estimation methods provide solutions in the form of probability distributions. Consequently, multiple potential solutions can be derived from the distributions, effectively addressing the inherent uncertainties within the data and the models relevant to the ECG inverse inference [19]. Specifically, the inverse problem can be formulated as the estimation of a Bayesian posterior probability distribution function (pdf), which can be defined as,
$p(\boldsymbol{u}|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{u},\epsilon)p(\boldsymbol{u}),$ (5)
where $p(\boldsymbol{u})$ is the prior, and the likelihood $p(\boldsymbol{y}|\boldsymbol{u},\epsilon)$ can be modeled as a normal distribution $\mathcal{N}(\boldsymbol{H}\boldsymbol{u},\epsilon\boldsymbol{I})$ [99]. One can predict the mode of the posterior distribution via maximum a posteriori (MAP) estimator, which is a point estimator that finds the parameter value that maximizes the posterior pdf given the observed data. When dealing with the complex dynamics of cardiac electrical activity, the Kalman filter serves as a specialized instance of the MAP estimator, enabling the integration of spatio-temporal information for the ECG inverse inference [100]. Alternatively, to estimate the entire posterior distribution, one can employ Markov Chain Monte Carlo sampling or variational inference for approximating analytically intractable posterior distributions. These approaches allow for a more comprehensive understanding of the cardiac electrical activity, taking into account the inherent variability and errors in the data and models, ultimately leading to more reliable and interpretable results.
2.3.1 Kalman Filter
The Kalman filter (KF) is a well-established tool for state estimation in dynamic systems. In the context of cardiac modelling, it implicitly estimates the posterior distribution of the underlying electrical activity given the observed ECG measurements. This is achieved by modelling the cardiac system as a high-dimensional stochastic state-space model, where the KF or its extensions can be effectively applied [79]. The key concept involves estimating the state $\boldsymbol{u}(\boldsymbol{s},t)$ by sequentially updating its estimate in the presence of noisy observations. The state transition is described as follows,
$\boldsymbol{u}(\boldsymbol{s},t)=\boldsymbol{F}\boldsymbol{u}(\boldsymbol{s},t-1)+\boldsymbol{w}_{t},$ (6)
where $\boldsymbol{F}$ refers to the state transition matrix (STM) and $\boldsymbol{w}_{t}$ is the process noise vector which models uncertainties and errors arising from the STM [101]. Then, the KF calculates the Kalman gain and uses it to update the state estimate based on observed data, considering spatio-temporal variations within cardiac source distributions. It provides not only a point estimate of the state but also a covariance matrix that represents the uncertainty associated with the estimate. By modelling both the dynamics of the system and the measurement noise probabilistically, KF inherently accounts for uncertainty in the estimation process.
To reduce computational complexity, researchers often focus on temporal relationships among spatially well-correlated regions, such as neighbouring regions or leads belonging to the same activation wavefront. They may also employ methods like expectation-maximization (EM) and residual-based techniques to estimate noise variances [101]. Other formulations, like the Duncan and Horn formulation of the KF used by Berrier et al. [71], focus on reconstructing specific electrograms. Wang et al. [79], in an attempt to integrate general prior knowledge with subject-specific data, used a statistical perspective to explicitly account for both model and data errors. They represented the cardiac electrophysiological system in a state-space form and conducted inverse inference of 3D TMP distributions via the unscented KF. A similar strategy was employed to define both the electrical and mechanical measurements as the state variables, which can be estimated via the unscented KF [102]. While the KF is a powerful tool for state estimation in cardiac modelling, it involves large matrices operation and may suffer from the initialization problem.
2.3.2 Markov Chain Monte Carlo Sampling
Figure 6: Example of the MCMM application on the inverse inference of cardiac activation properties from ECG. Here, MCMC has been used for the sampling of parameter sets of the population [103]. Image adapted from Camps et al. [103] with permission.
Current probabilistic inference approaches are generally based on Markov Chain Monte Carlo (MCMC) sampling methods to handle complex probabilistic models and uncertainty associated with ECG data [104]. For instance, Camps et al. [103] presented a sequential Monte Carlo approximate Bayesian computation-based model for the inference of ventricular activation properties. Rahimi et al. [105] applied a hierarchical Bayesian inference to estimate the statistical parameters in the posterior pdf, which can be calculated via MCMM sampling methods. However, direct MCMC sampling of the pdf of the parameters can be impractical, as it involves repeated evaluations of the posterior pdf which consists of computationally expensive simulation model. To accelerate sampling, various hybrid sampling techniques have been developed using information about the target pdf such as gradient and Hessian matrix [106], which however are challenging to extract when dealing with complex simulation models. Alternatively, one could develop computationally efficient surrogate models of the expensive simulation process, making it significantly faster to sample the associated pdfs [107, 108]. Despite improved efficiency, direct sampling of surrogate-based posterior pdf may yield limited accuracy, given the challenge of creating accurate approximations for complex nonlinear simulation models. Instead of completely replacing the sampling of the exact posterior pdf, Dhamala et al. [83] proposed a two-stage model to integrate Gaussian process surrogate modelling of the posterior pdf in accelerating sampling. Also, one could directly develop a surrogate model for the posterior pdf of simulation model parameters, eliminating the necessity for additional MCMC sampling of the computationally intensive original pdf [104].
2.3.3 Variational Inference
Variational inference (VI) approximates the posterior distribution with a simpler variational distribution by solving an optimization problem. This optimization process is often computationally more efficient than MCMC sampling as it does not require repeated sampling from the posterior distribution. Therefore, the application of VI has been widely adopted in solving the ECG inverse problem [74]. For instance, Xu et al. [99] introduced a Bayesian inference framework for robust transmural electrophysiological imaging. Ghimire et al. [74] proposed a Bayesian model for the estimation of the prior model error while reconstructing TMP under the constraint of a 3D EP simulation model. These conventional VI methods are generally optimized using techniques such as gradient descent and expectation-maximization (EM) algorithms.
Recently, deep learning-based probabilistic methods have served as an emerging alternative to conventional methods for modelling complex dynamics of cardiac electrical activity [109, 110, 111]. They can leverage deep neural networks to approximate the posterior distribution of the model parameters or latent variables, providing faster and more accurate approximations. For example, Ghimire et al. [109] proposed a deep generative model to reconstruct cardiac transmembrane potential from ECG data. Bacoyannis et al. [112] employed a variational auto-encoder (VAE) model to reconstruct the activation pattern of the myocardium with various local wall thicknesses as thin walls indicated infarct areas. Bacoyannis et al. [113] employed conditional VAE to generate activation maps from combined BSP and conductivity maps. Li et al. [114] designed a VAE-based deep computational model for the inverse inference of ventricular activation properties, i.e., root node (RN) position and CVs. Dhamala et al. [115] employed a graph convolutional VAE to allow generative modelling of non-Euclidean data for cardiac model personalization. Dhamala et al. [110] proposed a generative VAE for parameter estimation of a personalized cardiac model. Yang et al. [111] proposed Bayesian PINN to solve linear or nonlinear PDEs with noisy data for both forward and inverse problems. They employed the Hamiltonian Monte Carlo or the variational inference for the estimation of the posterior distributions.
Figure 7: Example of the conditional-VAE model for the inverse inference of myocardial infarction (MI) [116]. Image adapted from Li et al. [116] with permission.
2.3.4 Surrogate Model for Bayesian Optimization
To accelerate the inverse inference of parameters, surrogate models, especially statistical emulation with Gaussian processes (GP), have been employed in cardiovascular fluid dynamics [117], the pulmonary circulatory system [118], ventricular mechanics [119], and LV cardiac dynamics [117]. It involves approximating the computationally intensive mathematical model (the simulator) with a computationally efficient statistical surrogate model (the emulator) [117]. This is achieved through a combination of extensive parallelization and GP regression, allowing for optimal utilization of computational resources prior to the patient arrival at the clinic. Here, GP regression is a non-parametric Bayesian approach used for regression tasks where the model learns from the observed data to make predictions. It does not assume a specific parametric form for the underlying function but instead models the function as a distribution over functions, which is determined by the observed data points. Therefore, when new data are accessible, the proxy objective function can be minimized at a low computational cost. This eliminates the need for additional computationally intensive simulations from the original mathematical model and thus accelerates the inference. Recently, the GP surrogate model started to be used in the ECG inverse inference by Pezzuto et al. [120], where the GP surrogate model with Bayesian optimization has been used for the inverse inference of RNs. These strategies offer a mid-way solution between pure sampling methods using mechanistic models and machine learning-based solutions as the variational inference. They retain the flexibility and interpretability of pure sampling methods while replacing the mechanistic model with a faster emulator, which can be machine learning-based.
3 Uncertainty Quantification for the ECG Inverse Problem
Solving the ECG inverse problem relies on many subject-specific measurements and modelling assumptions, which introduce various sources of uncertainty in the predictions. Uncertainty quantification aims to analyze how models respond to inherent variability in input parameters. ECG inverse problems are often ill-posed and lack uniqueness in their solutions. However, clinicians typically require a single prediction for practical decision-making purposes in patient care scenarios. Therefore, it is essential to quantify and mitigate the effects of different types of uncertainties to ensure the robustness and clinical applicability of ECG inverse solutions.
3.1 Measurement Acquisition Uncertainty
ECG inverse inference requires subject-specific EP measurements, which can be easily distorted by noise artifacts, such as motion artifacts. EP data typically exhibit sparsity, with measured values associated with a higher degree of uncertainty [121]. For example, noise, a ubiquitous challenge in real-world measurements, can significantly distort the recorded ECG signals, introducing uncertainty into the source localization. To reduce the effects of measurement noise on the inverse problem, Aydin et al. [101] used a Kalman filter-based approach. Dogrusoz et al. [122] employed Bayesian estimation for solving the inverse ECG problem by applying various prefiltering methods to reduce the effects of noise. Similarly, Neumann et al. [123] utilized stochastic parameter estimation, aggregating probabilities estimated under various noise levels to derive a robust parameter estimate without explicitly determining the level of noise in the measurement. Moreover, there is ambiguity in matching the measurement points in the BSPM recording with the computational mesh derived from imaging data due to the non-synchronized BSPM-imaging acquisition [107]. The heart position is also susceptible to various sources of movement artifacts, including respiration and cardiac contraction, introducing geometric errors into the inverse problem of ECG solutions [124]. Therefore, the geometric error in the anatomical modelling (forward model) should also be considered, as it will in turn cause estimation errors in the ECG inverse inference.
3.2 Anatomical Modeling Uncertainty
The anatomical modelling is normally started with image segmentation, and then numerical solution techniques will be used to solve the forward ECG problem [58]. Therefore, the precision of the anatomical model employed, hinging on segmentation accuracy [125, 126], inter-subject anatomical variations [127], structural remodelling of the myocardium [116], and mesh resolution [128], profoundly influences the reliability of inverse solutions. Incorporating a prior model can mitigate uncertainties, and a combination of prior models may be preferred for estimating complex source structures [105]. Also, one can disentangle the inter-subject anatomical variation to mitigate its effect on the ECG inverse problem [127]. Furthermore, biventricular pacing and scar size, often encountered in clinical scenarios, demand meticulous modelling for a faithful representation [129]. Besides the heart, the torso geometry accuracy can also subtly influence the inverse solution, as the EP measurement is generally performed on the body surface [130]. In most studies, each torso surface node often corresponds to an electrode, and thus insufficient electrodes may lead to non-representative mesh nodes, requiring interpolation techniques for solving the inverse problem [131]. Dogrusoz et al. [131] investigated the effects of interpolation on the inverse problem of ECG. There also exist many studies exploring the effects of cardiac position inside the torso on the inverse inference [132, 133]. To address uncertainties, advanced regularization techniques that encompass prior knowledge, noise reduction strategies, and advanced data assimilation methods show promise [58, 134]. As the field advances, careful consideration of these effects becomes paramount for robust and clinically meaningful results in solving the ECG inverse problem.
Figure 8: The cardiac digital twin generation workflow combining cardiac imaging and ECG data, with ECG inverse inference for personalization. Image modified from Li et al. [116] with permission.
3.3 Computational Modeling Assumption Uncertainties
Model assumptions are necessary simplifications or idealizations of reality that are made to facilitate mathematical or computational modelling. These assumptions may include assumptions about tissue properties, boundary conditions, geometry, and physiological processes. In ECG inverse inference, assumptions are often made about the homogeneity of tissue properties, the behavior of electrical signals within the heart, and the relationship between measured ECG signals and underlying electrical activity [126, 135, 136]. While model assumptions are essential for simplifying complex systems, they can also introduce uncertainties and errors into the modelling process. For example, assuming uniform tissue properties throughout the heart and torso may lead to inaccuracies in the predicted electrical activity, especially in regions with significant heterogeneity [130, 53]. Similarly, assuming specific boundary conditions or physiological behaviors may not fully capture the complexity of real-world scenarios, leading to discrepancies between model predictions and observed data. ECG inverse inference also depends heavily on the assumptions on the initial distribution of heart electrical potential [135]. Determining the initial potential distribution at the onset of the cardiac cycle is particularly challenging due to the difficulty in accurately pinpointing the anatomical site of the earliest excitation in the heart, often associated with the intricate Purkinje networks [95]. Therefore, sophisticated modelling strategies are required to adequately account for the influence of these networks on ECG inverse inference [137, 138]. Once uncertainties are quantified, strategies can be employed to mitigate their impact on model predictions. The process of balancing model assumptions and uncertainty is often iterative, requiring continuous refinement and validation of models against experimental or clinical data. This iterative refinement process helps researchers to build more robust and reliable models over time.
4 Dataset and Evaluation
4.1 Dataset
The solution to the ECG inverse inference problem is typically implemented within a cardiac digital twin platform, as shown in Fig. 8. Therefore, for evaluation purposes, access to both anatomical data from cardiac imaging and electrical activity recording data from ECG or BSP mapping is necessary. While there are many public cardiac imaging data [139, 140, 141, 142, 143, 144] and ECG recordings [145, 146, 147, 148], the evaluation of ECG inverse inference methods requires paired imaging and ECG/ BSP mapping data from the same patient, which is less publicly available. UK Biobank (UKB) [149] and China Kadoorie Biobank (CKB) [150, 151] provide paired imaging and ECG dataset. MyoFit46 includes 500 participants of approximately 75 years+ to undertake high-density surface ECGI and cardiac MRI [152]. Strocchi et al. [153] provided a publicly available virtual cohort of twenty-four four-chamber hearts built from heart failure patients. Experimental Data and Geometric Analysis Repository (EDGAR) is another public dataset for the application and validation of ECGI techniques [154]. It includes paired BSPM and endocardial/ epicardial recordings from both animals (dog, pig, etc) and human and torso tank datasets. Additionally, information regarding the cardiac position within the torso is essential, often requiring torso image data. Several studies have attempted to reconstruct the torso directly from standard cardiac imaging for the localization of electrodes, which was challenging due to partial torso shape information [155, 156]. There are also many validation studies conducted in ex vivo torso tanks [157, 158, 159], in vivo large animal models [160, 161, 93], or in patients [162, 25]. Fig. 9 and Fig. 10 provide the two most commonly used ex vivo torso tanks.
Figure 9: Illustration of Utah Torso Tank, with an isolated canine heart. Image adapted from Bergquist et al. [157] with permission.
Figure 10: Illustration of Bordeaux Torso Tank, with a pig heart perfused in Langendorff mode with a 108-electrode sock placed over the ventricles. Image adapted from Bear et al. [158] with permission.
Figure 11: Illustration of different cardiac source models in ECG.
4.2 Evaluation
The validation of the inverse problem involves a diverse set of cardiac source models, forward EP model formulations, and inverse methods, as presented in Table 3. For simplicity, we organized the validation according to the estimated cardiac source models, including surface potential models, activation/ recovery models, and transmembrane voltage models [52]. Fig. 11 provides three representative examples corresponding to the three source models. Except for a few studies, all the others employed BSP as the input representation/ measurement for cardiac electrical activity assessment. One can see that most studies employed computer models for in silico evaluation [66, 163] and/ or performed evaluation on the basis of torso-tank experiments with isolated animal hearts, such as rabbit [164], canine [77, 73, 92], and swine [165, 166, 167]. There are also few studies evaluated using in-vivo human subjects, which however usually employed non-simultaneous invasive recordings, such as EGMs [168, 169] or bipolar voltages maps [81, 74]. The comparison with the published invasive measurements in animals or humans has also been utilized as an alternative evaluation [24].
Table 3: Summary of the ECG inference works. EP: electrophysiology; AS: activation sequence; EGM: electrograms; HF: heart failure; HS: healthy subject; EpiP/ EndoP: epicardial/ endocardial potential; RN: root node; VCG: vectorcardiographic; AP: Aliev-Panfilov; LSTM: long short-term memory; FCN: fully connected neural network; TV: total-variation.
Source & Publish year Data Input Target EP model Inverse inference method
Ramanathan et al. (2004) [24] 1 HS; 3 patients BSP EpiP, EGM, isochrones boundary element Tikhonov regularization + generalized minimal residual algorithm
Shou et al. (2008) [60] in silico BSP EpiP finite element truncated total least squares
Wang et al. (2009) [63] in silico BSP EpiP finite element Tikhonov regularization
Jiang et al. (2009) [170] in silico BSP EpiP boundary element subspace preconditioned LSQR
Sarikaya et al. (2010) [59] 1 canine BSP EpiP N/A genetic algorithm with conventional regularization
Wang et al. (2010) [171] 1 canine BSP EpiP finite element variational-form-based regularizers
Wang et al. (2011) [172] 1 canine BSP EpiP finite element variational regularizers
Cluitmans et al. (2013) [168] 1 patient BSP EpiP N/A realistic data and VCG based optimization
Zemzemi et al. (2013) [173] in silico BSP EpiP bidomain Steklov-Poincare formulation + gradient descent method
Zemzemi et al. (2014) [174] in silico BSP EpiP bidomain iterative Kozlov-Maz’ya-Fomin
Erem et al. (2014) [82] in silico; 1 HS BSP EpiP, EndoP N/A temporal spline model
Yao et al. (2017) [175] 4 MI patients BSP EGM, MI N/A spatiotemporal regularization
Cluitmans et al. (2017) [176] 3 canine BSP EpiP AP physiology-based regularization
Cluitmans et al. (2018) [73] 3 canine BSP EpiP, ATM, beat origin N/A multitask elastic-net-based method
Kalinin et al. (2019) [163] in silico BSP EpiP, EndoP bidomain single layer approach
Karoui et al. (2019) [88] in silico BSP EpiP monodomain spatial adaptation of time delay neural network
Bear et al. (2020) [65] 4 canine/swine BSP EpiP, ATM boundary element Tikhonov regularization
Rababah et al. (2021) [67] 6 canine/swine BSP EpiP, ATM boundary element Tikhonov regularization
Ozkoc et al. (2021) [177] 1 canine BSP EpiP boundary element Bayesian MAP
Chen et al. (2022) [167] 5 swines ECG EpiP, ATM, RN N/A FCN, CNN, LSTM model
Tenderini et al. (2022) [85] in silico ECG EpiP AP PDE-aware deep learning model
Jiang et al. (2022) [92] 1 canine BSP EpiP AP physics embedded graph CNN
Xie et al. (2023) [178] in silico BSP EpiP AP physics-constrained network
Jiang et al. (2024) [179] in silico; 3 patients BSP EpiP, EndoP reaction-Eikonal model neural state-space modeling
Messnarz et al. (2004) [68] in silico BSP TMP bidomain spatiotemporal regularization
Wang et al. (2009) [79] in silico BSP TMP bidomain physiological-model-constrained statistical approach
Wang et al. (2011) [180] in silico BSP TMP, MI bidomain physiologically-based constraints
Potyagaylo et al. (2012) [181] in silico BSP TMP N/A iterative Gauss-Newton optimization
Jiang et al. (2013) [182] in silico BSP TMP bidomain Support vector regression
Lopez et al. (2014) [183] in silico BSP TMP monodomain genetic algorithm with Tikhonov regularization
Xu et al. (2014) [81] in silico; 6 patients BSP MI N/A variational Bayesian with TV prior
Zettinig et al. (2014) [38] in silico ECG Myo electrical diffusivity LB polynomial regression
Ghimire et al. (2019) [74] in silico; 2 patients ECG TMP finite element Bayesian model
Kara et al. (2019) [64] in silico BSP TMP, MI bidomain Tikhonov regularization
Cheng et al. (2021) [184] in silico BSP TMP, MI bidomain shrinkage-thresholding network
Mu et al. (2021) [185] in silico BSP TMP N/A shrinkage-thresholding + graph CNN
Zhao et al. (2022) [186] in silico BSP TMP bidomain data-driven Kalman filtering network
Li et al. (2024) [116] in silico ECG MI Eikonal multi-modal VAE
Lian et al. (2024) [187] in silico; 4 patients ECG MI bidomain frequency-enhanced model
Zhang et al. (2005) [164] 4 rabbits BSP ATM N/A N/A
Calderero et al. (2005) [77] 1 canine BSP activation wavefront potential-based level sets
Alawad et al. (2019) [188] in silico; 3 patients ECG pacing sites N/A L1-norm regularization
Schuler et al. (2021) [66] in silico BSP ATM bidomain Tikhonov regularization
Pezzuto et al. (2021) [169] 11 HF patients ECG ATM, RN Eikonal propagation model
Meister et al. (2021) [165] 16 swine ECG ATM N/A graph convolutional regression
Camps et al. (2021) [103] in silico ECG CV, RN N/A Monte Carlo Bayesian inf
Li et al. (2022) [189] in silico ECG CV, RN Eikonal conditional VAE
Camps et al. (2024) [138] in silico ECG CV, Purkinje network Eikonal Monte Carlo Bayesian inf
4.2.1 Surface Potential Model
The prevalence of potential map estimation on the cardiac surface can be attributed to its practicality and clinical significance. It provides a powerful and accessible means of understanding cardiac electrical behavior. Electrophysiological information reconstructed on the cardiac surfaces, such as the epicardium or endocardium, is relatively easier to obtain and is invaluable for clinical insights. Most surface potential inference work primary focused on the reconstruction of epicardial (outer) potential rather than endocardial (inner) potential. This is reasonable, as epicardial potential is more accessible for measurement compared to endocardial potential. As for the forward EP model, boundary element, finite element, and Aliev-Panfilov (AP) model have been widely used.
4.2.2 Activation/ Recovery Model
Activation and recovery time maps offer valuable static isochrone images depicting the sequence of cardiac electrical depolarization and repolarization. These isochrones can be applied in the clinic in various aspects: (a) identifying sites where activation initiates, as seen in focal arrhythmias, e.g., premature ventricular contractions (PVC) origin; (b) mapping reentry circuits, e.g., in macroreentrant atrial flutter; (c) detecting abnormalities in propagation, including areas with delayed activation [190] or those exhibiting slow or abnormal conduction, e.g., scar-related ventricular tachycardia (VT) substrate; and (d) pinpointing areas with recovery abnormalities, e.g., large recovery gradients [191, 192]. To date, the majority of validation studies have concentrated on the accuracy of activation mapping, with only a few studies assessing recovery [76, 193].
Besides, one could estimate activation parameters for personalization. The estimation of EP parameters from ECG or BSP is relatively rare, as these parameters are typically more detailed and are best captured through invasive electrophysiological studies or high-resolution mapping techniques. For example, Grandits et al. [194] employed Huber regularization to estimate simultaneously distributed fibre architecture and CVs from sparse contact recordings.
4.2.3 Transmembrane Potential Model
Using the transmembrane potential (TMP) as a cardiac source model offers a highly accurate approximation of the cardiac actual electrical source [52]. Estimating volumetric myocardial TMP dynamics is clinically significant as it unveils intricate arrhythmic details and intramural arrhythmogenic substrates. However, this necessitates complex modelling and interpretation of electrical activity within the 3D cardiac structure. Reconstructions of TMP support the identification of depolarized and repolarized regions in the heart and their temporal evolution. They also delineate areas of reduced amplitude or completely absent TMP, such as ischemic, fibrotic, or border-zone areas. Nevertheless, TMPs cannot be directly measured in the clinic, necessitating indirect validation via extracellular signals generated by TMP distributions. Surface potentials on the endocardial and epicardial surfaces, as well as activation and recovery times, can be calculated from TMP distributions and then measured in experiments.
5 Clinical Application
Inverse modelling of electrocardiographic signals can help to create a personalized virtual heart, which can be potentially applied for: (a) preoperative planning for cardiac surgery, (b) risk stratification and long-term monitoring, and (c) personalized disease insight and diagnosis.
Figure 12: Virtual heart model guided preoperative planning for the personzalized catheter ablation of atrial fibrillation (AF) patients. Here, CARTO is the clinical electroanatomic navigation system. Illustrations designed referring to Boyle et al. [3].
Figure 13: Comparison of conventional clinical metric and virtual heart model based identification of patients at high sudden cardiac death (SCD) risk. (a) a universal metric in clinical cardiology practice, (b) a personalized virtual-heart arrhythmia risk predictor. ICD: implantable cardioverter defibrillator. Illustrations designed referring to Arevalo et al. [2].
5.1 Preoperative Planning for Cardiac Surgery
Virtual heart models enable surgeons to simulate different surgical interventions and assess their feasibility and potential outcomes in a risk-free environment. For instance, utilizing personalized virtual-heart technology enables the identification of optimal infarct-related ablation targets for VT [193, 195, 196]. This can be accomplished by performing virtual multi-site delivery of electrical stimuli (pacing) from various bi-ventricular locations. Each location attempts to induce VT from a site positioned differently relative to the infarct area. The optimal ablation lesions can then be determined in each personalized virtual heart, rendering it non-inducible for VT from any pacing location. Therefore, it can terminate not only clinically manifested or induced VTs during the procedure but also all potential VTs arising from the post-infarction substrate. This includes VTs that may emerge after the initial ablation, potentially eliminating the need for repeated procedures and providing long-term freedom from VT for patients. Similarly, this technology allows for the preoperative identification of optimal ablation targets for AF, as shown in Fig. 12. It has proved invaluable in guiding ablation strategies, enabling clinicians to target specific areas responsible for atrial arrhythmias more effectively [3]. For MI management, ECG inverse inference aids in delineating the extent and location of ischemic regions within the myocardium [74, 116]. It can also be utilized to detect pulmonary vein (PV) drivers and thus assist in identifying patients with high acute success rates to PV isolation [197]. Furthermore, surgeons can virtually implant cardiac devices such as pacemakers, defibrillators, or transcatheter valves to evaluate their placement and function. Additionally, complex surgical procedures such as coronary artery bypass grafting and valve repair or replacement can be simulated to optimize surgical strategies, minimize operative risks, and improve patient outcomes.
5.2 Risk Stratification and Long-Term Monitoring
In risk stratification, ECG inverse inference techniques contribute by unravelling intricate electrophysiological patterns that signify an increased risk of adverse cardiac events. For instance, Arevalo et al. [2] created a virtual-heart arrhythmia risk predictor (VARP) for sudden cardiac death (SCD) based on a virtual heart model, as shown in Fig. 13. The heart model integrates the ventricular geometry of the patient, structural remodelling resulting from post-MI and electrical functions from sub-cellular to organ levels. Within each heart model, a virtual multi-site delivery of electrical stimuli originating from ventricular locations at varying distances to the remodelled tissue can be performed. This comprehensive evaluation allows for a thorough assessment of the patient’s heart propensity to develop ventricular arrhythmias related to MI. The VARP has been demonstrated to significantly outperform clinical metrics in predicting future arrhythmic events. The resilient and non-intrusive VARP approach holds promise in potentially averting SCD and steering clear of unnecessary implantable cardioverter-defibrillator (ICD) placements in patients with post-MI. This is particularly valuable in monitoring patients with known cardiovascular conditions, where early signs of deterioration can be identified, enabling timely intervention and personalized adjustments to the treatment plan [198, 199]. This can be achieved by combining multiple sources (fitness trackers, wearables, and ’companion’ apps for drugs, etc.) to maintain and update cardiac digital twins and their underlying models. The integration of ECG inverse inference thus emerges as a powerful tool in risk assessment and continuous monitoring, contributing to more proactive and personalized cardiovascular care.
5.3 Personalized Disease Insight and Diagnosis
Numerous studies within the pathological validation section have delved into disease mechanisms using ECG inverse inference, proposing clinically relevant parameters [200, 201, 202]. One of the earliest extensively clinically explored applications of ECG inverse inference is in identifying the specific mechanisms underlying arrhythmia onset and/ or perpetuation for ablation therapy [52]. For instance, it was utilized to delineate the distinct activation patterns during arrhythmia (e.g., focal vs. re-entrant) [203, 204]. Additionally, it has also been employed to pinpoint the drivers of this activation [205], which have demonstrated correlation with successfully ablated regions in patients with positive outcomes and are now the focus of direct targeting in clinical validation studies [206]. Besides atria, ECG inverse inference has also been applied to identify and understand arrhythmogenic substrates for the onset and maintenance of VT [201, 207] and ventricular arrhythmias [208]. It has showcased larger ventricular electrical dyssynchrony in CRT responders, improving patient selection over standard 12-lead ECG [209, 210]. ECG inverse inference techniques are especially valuable in scenarios where invasive recordings are ethically constrained, such as with infants and pregnant women, offering insights into cardiac evolution. They can also be used for in silico clinical trials (see Section 6.3), which can particularly facilitate safe methods to explore treatment effects in patients with rare diseases and thus may allow for insights not possible in the current clinical trial practice [35]. Certain diagnostic medical devices, like CardioInsight (Medtronic, USA), based on personalized models, have transitioned into industrial application and clinical usage [35]. Within the CardioInsight mapping system, the cardiac electrical activity on its surface can be reconstructed from body surface potentials employing a customized model of the heart and torso of the target patient [211]. While these retrospective studies contribute to understanding disease mechanisms and defining potential therapeutic markers, further exploration through extensive prospective studies is crucial to establish the clinical value of these insights for individual patients and their potential role in guiding diagnosis or patient selection.
6 Discussion and Future Perspectives
6.1 Surrogate Model for Efficient ECG Simulation
Traditional methods for simulating ECG signals involve complex biophysical models that account for intricate cardiac electrical and anatomical properties. While these models offer high fidelity, they often require substantial computational resources and time [5]. Fast implementations of these models exist. For example, Sachetto et al. [212] presents a GPU-based implementation of the monodomain model, which is capable of speeding up simulations by a factor up to 498 when compared to a serial implementation of the model. They reported simulation times of 32 minutes per heartbeat. Other studies have proposed alternative models that can yield similar simulation outputs at a fraction of the state-of-the-art monodomain models. For example, Wallman et al. [121] proposed a graph-based Eikonal model that can simulate the activation sequence in the heart in less than 1.6 seconds [103]. Similarly, Neic et al. [213] presented a pseudo-monodomain model that can produce nearly identical simulations to a monodomain model (under sinus rhythm) but with a computational cost of 213 seconds per heartbeat [214]. Nevertheless, in the context of generating digital twins, the number of simulations required per patient can easily be in the range of hundreds of thousands [103, 138]. Thus, even the fastest mechanistic models available present a significant computational burden when designing studies involving the generation of digital twins at scale [214, 138].
The emergence of neural surrogate models addresses these challenges by offering computationally efficient alternatives that balance accuracy with practicality [215, 216, 217]. These surrogate models leverage machine learning techniques to learn the intricate mapping between the physiological parameters (e.g., earliest activation sites) and expected behavior (e.g., cardiac activation sequences) [217]. For example, Bertrand et al. [217] was able to produce simulations of the activation sequence in 0.04 seconds, which was a 500-factor speed-up compared to a serial implementation of the graph-based Eikonal [121] (20 seconds per simulation). By circumventing the need for exhaustive simulations, surrogate models provide an avenue for rapid and on-demand generation of cardiac functional data (e.g., ECG signals). This is particularly valuable in scenarios where real-time simulations or large-scale analyses are necessary. However, challenges such as model interpretability, generalization across diverse populations, and the need for accurate training data warrant careful consideration. Furthermore, a thorough understanding of the trade-off between computational efficiency and simulation accuracy is vital when integrating surrogate models into clinical applications. As surrogate models continue to evolve, they hold the potential to revolutionize ECG inverse inference by enabling researchers and practitioners to efficiently explore the parameter space and improve our understanding of complex cardiac phenomena.
6.2 Multimodal Representation Learning Towards Solving the ECG Inverse Problem
The ECG inverse inference involves the fusion of information from multiple modalities, such as ECG signals, anatomical images (e.g., CT/ MRI scans), and clinical metadata. Multimodel representation learning offers a promising solution by allowing the integration of diverse data sources, leading to improved insights into the complex dynamics of cardiac electrical behavior. This can help address the issue of limited electrode coverage by inferring missing data points through learned relationships from other modalities. Furthermore, the inclusion of anatomical images can enhance the anatomical accuracy of the models, leading to improved source localization and more accurate inverse solutions. Indeed, deep neural networks have proven to be a highly effective strategy for ECG interpretation and analysis via spatiotemporal representation learning [47]. However, most works only considered ECG data for analysis and only a few works employed joint analysis with other data sources [218, 189, 219, 116]. Among these works, VAE has been widely employed to learn a latent representation of ECG and geometry [220], along with incorporating uncertainty modelling for inverse inference [116]. However, these studies merely concatenated the acquired features or utilized a shared representation, lacking a thorough explanation of how these features align and contribute to holistic data comprehension. In contrast, physics-informed models can integrate domain-specific knowledge and principles derived from the underlying physics of the cardiac system [179]. By explicitly encoding the physical interactions and dependencies between ECG and geometry, these models could offer a more transparent and interpretable way to capture the intricate interplay between these data modalities. Moreover, they could potentially enhance the generalization capabilities of the learned representations, making them more robust in scenarios with limited or noisy data. Nevertheless, the integration of physics-based constraints into machine learning models introduces additional challenges, such as learning geometrical invariance, as well as processing model complexity and the need for accurate domain knowledge. Balancing the complexity of the physical model with the capacity of the machine learning algorithm while also ensuring the interpretability of the resulting features will require careful consideration and innovative methodologies.
6.3 The Role of Cardiac Digital Twins in Advancing in Silico Clinical Trials
Cardiac digital twins are revolutionizing the landscape of in silico clinical trials, offering unprecedented opportunities to accelerate drug development [138]. By creating virtual replicas of the individual patient’s heart, cardiac digital twins enable highly personalized simulations that capture the intricate dynamics of cardiac function and response to interventions [221]. Traditional clinical trials often face limitations in recruiting participants from specific demographic or clinical groups, leading to challenges in generalizing findings to broader populations [222]. With cardiac digital twins, researchers can create virtual patient cohorts that span a wide range of demographic, clinical, and pathological characteristics. This diversity allows for a more comprehensive exploration of drug effects and treatment outcomes across different patient subgroups, enhancing the reliability and generalizability of trial results. At the same time, in silico trials significantly reduce the time and cost associated with recruiting participants, conducting experiments, and analyzing results. This accelerated pace facilitates the rapid iteration and refinement of drug candidates, ultimately expediting the journey from discovery to market availability. Moreover, in silico trials enable researchers to explore a broader range of hypotheses and scenarios than would be feasible in traditional clinical settings. This flexibility allows for a more comprehensive exploration of potential drug effects, interactions, and mechanisms of action, leading to deeper insights and more informed decision-making. In the future, in silico clinical trials hold immense potential for revolutionizing drug development and healthcare delivery. The integration of patient-specific data and virtual patient models will enable personalized drug development approaches tailored to individual characteristics and needs. To fully realize the potential of in silico clinical trials, challenges such as model validity, regulatory compliance, and data access barriers must be addressed.
7 Conclusion
Cardiac digital twins are pivotal in advancing precision medicine and improving patient outcomes in cardiovascular care by simulating and predicting personalized responses to interventions. However, current models encounter a computational bottleneck in calibrating to an individual’s physiology through ECG inverse inference, particularly in 3D heart models. This review comprehensively outlines recent progress in ECG inverse inference, encompassing algorithms, public datasets, evaluation metrics, clinical applications, and future perspectives. While machine learning, especially deep learning, holds promise in tackling this complex and ill-posed problem, significant challenges persist, alongside promising opportunities for further advancement and innovation. In the future, as these techniques gain increasing popularity and generate massive cohorts of digital twins, they are poised to revolutionize cardiovascular care through in-silico trials and personalized interventions. With the proliferation of big data, envision a future where everyone possesses a digital twin, significantly reducing the reliance on animal models and enabling real-time monitoring in hospitals.

### v4

@@
 Abstract
-Cardiac digital twins are personalized virtual representations used to understand complex heart mechanisms.
-Solving the ECG inverse problem is crucial for accurate virtual heart modelling, enabling the derivation of internal electrical activity information from recorded surface potentials.
-Despite challenges from cardiac complexity, noisy ECG data, and computational efficiency, recent advancements hold significant promise for enhancing virtual heart modelling, ultimately advancing precision medicine in cardiology.
+Cardiac digital twins (CDTs) are personalized virtual representations used to understand complex cardiac mechanisms.
+A critical component of CDT development is solving the ECG inverse problem, which enables the reconstruction of cardiac sources and the estimation of patient-specific electrophysiology (EP) parameters from surface ECG data.
+Despite challenges from complex cardiac anatomy, noisy ECG data, and the ill-posed nature of the inverse problem, recent advances in computational methods have greatly improved the accuracy and efficiency of ECG inverse inference, strengthening the fidelity of CDTs.
 This paper aims to provide a comprehensive review of the methods of solving ECG inverse problem, the validation strategies, the clinical applications, and future perspectives.
-For the computing methodologies, we broadly classify state-of-the-art approaches into two categories: deterministic and probabilistic methods, including conventional and deep learning-based techniques.
+For the methodologies, we broadly classify state-of-the-art approaches into two categories: deterministic and probabilistic methods, including both conventional and deep learning-based techniques.
 Integrating physics laws with deep learning models holds promise, but challenges such as capturing dynamic electrophysiology accurately, accessing accurate domain knowledge, and quantifying prediction uncertainty persist.
 Integrating models into clinical workflows while ensuring interpretability and usability for healthcare professionals is essential.
-Overcoming these challenges will drive further research in cardiac digital twins.
+Overcoming these challenges will drive further research in CDTs.
 {IEEEkeywords}
-Cardiac electrophysiology, inverse problem, computational model, cardiac digital twin, survey.
+Cardiac digital twins, ECG, inverse problem, ECGI, parameter estimation, model personalization, survey.
 1 Introduction
-Electrocardiogram (ECG) is a non-invasive technique used to measure the electrical activity of the heart and identify potential abnormalities [1].
-It consists of a number of electrodes placed on the skin of the patient, typically on the chest, arms, and legs.
-These electrodes record the changes in electrical potential on the body surface, which are generated by the underlying electrical activity of the heart.
-Virtual modelling of cardiac electrophysiology (EP) for digital twins has shown great promise in numerous clinical applications, such as the stratification of arrhythmia risk [2] and the ablation guidance of persistent atrial fibrillation (AF) [3].
-However, obtaining numerical solutions to the governing partial differential equations (PDEs) in cardiac EP (ECG forward problem) is quite challenging.
-Several forward physical modelling approaches have been employed, such as the bidomain model, the monodomain model, and the Eikonal equation.
-The biodomain model provides information about the ionic currents [4], the monodomain model describes the electrical propagation [5], and the Eikonal equation models wavefront propagation [6].
-To obtain a complete understanding of the electrical activity of the heart, we need to estimate the activity inside the heart from the measured ECG signals, namely ECG inverse problem [7].
-Fig.
-1 presents the inverse inference procedure for reconstructing the cardiac activity from ECG data.
-By solving the inverse problem, one can obtain valuable information about the function of the heart, which can aid in the diagnosis and monitoring of various cardiac conditions, as well as the development of personalized treatments for cardiac disorders.
-However, directly predicting cardiac activity from standard ECG signals (often recorded using a limited number of electrodes, such as the 12-lead ECG with ten electrodes) is a challenging task.
-This is mainly because the limited number of electrodes in standard ECGs makes it difficult to capture the fine-grained details of cardiac electrodynamics across the entire heart.
-Therefore, body surface potential (BSP) mapping has been developed by deploying a larger number of ECG electrodes, such as Dalhousie Standard Torson 120 electrodes, to provide high-resolution spatiotemporal information about cardiac electrodynamics projected onto the body surface [8].
-It allows for detailed mapping and localization of cardiac activity and thus has been widely used in the last decades for solving the ECG inverse problem.
-Nevertheless, the ECG inverse problem, like many other inverse problems, faces two primary challenges.
-Firstly, the problem is ill-posed, meaning that it may not have a unique solution and small errors in the input data can lead to significant errors in the output.
-Secondly, solving the ECG inverse problem can be computationally intensive, particularly when dealing with high-resolution BSP mapping data.
-The complex geometrical, mechanical, and electrical properties of the heart, as well as the need for real-time or near-real-time solutions in clinical applications, pose computational challenges.
-Figure 1: Illustration of the electrocardiogram (ECG) inverse problem.
-Table 1: Search engines and expressions used to identify potential papers for review.
-Engine Google Scholar, PubMed, IEEE-Xplore, and Citeseer
-Term “Electrocardiogram” or “ECG” or “EKG” or “electrocardiographic” or “electrophysiological” or “electrophysiology” or “EP” or “electrocardiography” or “body surface potential mapping” or “BSPM” and
-“Inverse problem/ issue” or “inverse solution/ estimation/ inference/ reconstruction” or “cardiac model personalization” or “electrocardiographic imaging/ mapping” or “ECG-imaging” or “ECGI” or “activation imaging” or “noninvasive imaging” and
-“Electrical activity/ excitation” or “surface potential” or “epicardial/ endocardial potential” or “action potential” or “transmembrane potential” or “activation sequence” or “activation wavefront” or “activation time map” or “pacing site” or “atrial potential” or “local parameter”
-Figure 2: Timeline of the development of ECG, imaging, and cardiac digital twins.
+Electrocardiography (ECG) investigates the relationship between the cardiac electrical activity and the resulting potential measured on the torso surface.
+Understanding this relationship is crucial for developing personalized virtual models of cardiac electrophysiology (EP), i.e., cardiac digital twins (CDTs) [1, 2].
+Currently, treatment decisions are still primarily based on empirical clinical studies that statistically compare the effects of different treatments across large groups of patients with similar cardiac conditions [3].
+Instead, the virtual heart models have shown great promise in providing personalized treatments, such as arrhythmia risk stratification [4] and ablation guidance for persistent atrial fibrillation (AF) [5].
+This involves using computational models to solve the ECG forward and inverse problems.
+The ECG forward problem aims to predict signal on the body surface from known cardiac activity.
+Instead, the ECG inverse problem seeks to infer cardiac information from ECG or body surface potential (BSP) [6, 2], as presented in Fig.
+1.
+It can be conducted in two aspects, i.e., the reconstruction of cardiac sources and the estimation of EP parameters.
+Both aspects are closely linked with the concept of CDTs [7, 2].
+The reconstruction of cardiac sources provides crucial information about the cardiac electrical activity, which can be integrated into digital twin models to enhance their accuracy in representing the heart [7].
+On the other hand, the estimation of EP model parameters allows these digital twins to be fine-tuned to reflect individual patient conditions, ensuring that the model behavior closely matches the real cardiac physiological responses [7].
+Figure 1: Illustration of the electrocardiogram (ECG) inverse problem, including non-invasive cardiac source reconstruction, also known as ECG imaging (ECGI), and electrophysiological (EP) parameter estimation.
+Here, we take the biventricular modeling as an example.
+BSP: body surface potential; CV: conduction velocity.
+However, due to the nature of the ECG inverse problem, the reconstruction of cardiac sources and the estimation of EP parameters share several common challenges.
+First, both scenarios involve solving an inherently ill-posed inverse problem, which may not have a unique solution.
+This means that small errors in the input data can lead to significant inaccuracies in the output.
+Second, ECG data recorded from the body surface, often using a limited number of electrodes such as the 12-lead ECG with ten electrodes, has limited spatial resolution and is susceptible to noise, artifacts, and inconsistencies.
+Third, both scenarios need to account for the complex, inhomogeneous structure of the thorax, including the heart, lungs, and other tissues, which influence how electrical signals propagate to the body surface.
+Accurately modeling these structures poses a significant challenge that affects the reliability of reconstructed cardiac sources and the accuracy of estimated EP parameters.
+Compared to the inverse inference of cardiac sources, the inference of EP parameters presents several additional challenges.
+While the cardiac source reconstruction requires computational resources, it is generally not considered excessively time-consuming.
+In contrast, EP parameter estimation is far more computationally intensive due to its reliance on sophisticated anatomical modeling, iterative simulations, and complex parameter fitting, particularly when using high-resolution models.
+Also, EP parameters can be interdependent, meaning that changes to one parameter may affect others, which further complicates the optimization and fitting process [8].
+Unlike standard ECG imaging, which is directly validated against real-time measurements, the EP parameter estimation requires comprehensive validation to ensure the generated CDTs accurately represent the unique anatomy and EP of the patient.
 1.1 History: From Electrocardiography to Cardiac Digital Twins
@@
 This laid the groundwork for cardiac simulations using computer models.
-Meanwhile, body surface potential mapping (BSPM) emerged as an alternative method for non-invasive and high-resolution measurement of human cardiac electrical activity in 1963 [16].
+Meanwhile, BSP mapping emerged as an alternative method for non-invasive and high-resolution measurement of human cardiac electrical activity in 1963 [16].
 In 1970, Durrer et al.
 [17] conducted a seminal investigation into the electrical behavior of the human heart at the organ level.
-Since 1972, numerous studies have tackled the electrocardiography inverse problem using analytical and computational models [18, 19, 20, 21], experimental animals [22, 23], and human subjects [24, 25].
-Advancements in medical imaging have also enabled non-invasive visualization of cardiac anatomy since the inception of techniques like the first human MRI scan in 1977 [26].
-The formalization of inverse procedures reconstructing bioelectric sources from BSPM and medical imaging led to the development of electrocardiographic imaging (ECGi) techniques, notably pioneered by Yoram Rudy’s lab [27, 28, 29, 30, 31].
+Since 1972, numerous studies have tackled the electrocardiography inverse problem using analytical and computational models [18, 19, 20, 21, 22], experimental animals [23, 24], and human subjects [25, 26].
+The formalization of inverse procedures reconstructing bioelectric sources from BSP mapping (BSPM) led to the development of non-invasive electrocardiographic imaging (ECGI), a term formally coined by Yoram Rudy’s lab in 1997 [27].
+While ECGI can be patient-specific in terms of geometry and electrical potentials, it does not typically include a full-scale, personalized model of the cardiac EP properties.
+Table 1: Search engines and expressions used to identify potential papers for review.
+Engine Google Scholar, PubMed, IEEE-Xplore, and Citeseer
+Term “Electrocardio*” or “ECG” or “EKG” or “electrophysiolog*” or “EP” or “body surface potential” or “BSP” and
+“Inverse problem/ issue” or “inverse solution/ estimat*/ infer*/ reconst*” or “electrocardiographic imaging/ mapping” or “ECG-imaging” or “ECGI” or “activation imaging” or “noninvasive imaging” or “personaliz*/ calibrat*” and
+“Electrical activity/ excitation” or “heart/ epicardial/ endocardial /atrial / transmembrane potential” or “activation sequence/ wavefront/ time” or “pacing site” or “parameter” or “onset location/ root note/ earliest activation site/ Purkinje” or “conducti*” or “origin of cardiac activation” or “excitability”
+In constrast, CDT is a comprehensive, personalized model that incorporates both the electrical and structural properties of the heart, including detailed EP parameters.
 The concept of the digital twin was formally introduced in 2002 by Michael Grieves, though it can be traced back to NASA’s use for the Apollo 13 rescue mission in 1970 [28].
-The Physiome Project was initiated in 2003 with the aim of establishing a framework for modelling the human body, utilizing computational methods to integrate biochemical, biophysical, and anatomical information across cells, tissues, and organs [33].
+The Physiome Project was initiated in 2003 with the aim of establishing a framework for modeling the human body, utilizing computational methods to integrate biochemical, biophysical, and anatomical information across cells, tissues, and organs [29].
 Starting from 2014, many companies like Dassault, Siemens, and ANSYS embraced the term “digital twins” in their marketing initiatives.
-For example, Dassault Systèmes initiated the Living Heart Project in 2014, which is the first comprehensive computer model of the human heart, integrating all functional aspects, i.e., blood flow dynamics, mechanical properties, and electrical conduction [34].
-In recent years, digital twin tools have become popular in healthcare, such as cardiac digital twins [35], brain digital twins [36], etc.
-In 2016, the US Food & Drug Administration (FDA) Center for Devices and Radiological Health released the first guidance to enable the use of modelling and simulation to develop in silico trials [37].
-In 2019, Trayanova’s lab received FDA approval for a randomized clinical trial involving 160 patients, termed OPTIMA (OPtimal Target Identification via Modeling of Arrhythmogenesis), to demonstrate the effectiveness of cardiac digital twins in guiding atrial ablation procedures [3].
-The brief timeline has been presented in Fig.
-2.
-1.2 Study Inclusion and Literature Search
-In this work, we aim to provide readers with a survey of the state-of-the-art solutions to electrocardiographic inverse problems.
-There exist several modalities for cardiac electrical activity assessment, such as ECG, BSP mapping, electrograms (EGM), electro-anatomic mapping (EAM), vectorcardiography (VCG), etc.
-However, in the context of solving the ECG inverse problem, we usually only involve ECG and BSP as the observed data.
-Note that compared to standard 12-lead ECG, BSP mapping is, however, not yet widely available as a diagnostic modality [38].
-For cardiac activity representation, two major source models are employed, i.e., heart surface potential (HSP) and transmembrance potential defined on the volumetric mesh of the heart.
-Specifically, it can be further categorized into epicardial and/ or endocardial potential, activation time map/ activation wavefront, transmembrane potential, action potential duration map, voltage potential maps, isochrones/ isochrone maps, etc.
-Table 2: Summary of the related representative existing surveys. DL: deep learning; MCG: Magnetocardiography; ATM: activation time maps; AF: atrial fibrillation.
-Source & Publish Year Scope Inclusion
-Macleod et al. (1998) [7] ECG inverse problem conventional solutions
-Dossel et al. (2000) [39] ECG and MCG inverse problem conventional solutions
-Yaman et al. (2013) [40] inverse problem for applied sciences not for ECG; conventional solutions
+For example, Dassault Systèmes initiated the Living Heart Project in 2014, which is the first comprehensive computer model of the human heart, integrating many functional aspects, i.e., blood flow dynamics, mechanical properties, and electrical conduction [30].
+In recent years, digital twin tools have become popular in healthcare, such as CDTs [31], brain digital twins [32], etc.
+In 2016, the US Food & Drug Administration (FDA) Center for Devices and Radiological Health released the first guidance to enable the use of modeling and simulation to develop in silico trials [33].
+Notably, Trayanova’s lab received FDA approval in 2019 for a randomized clinical trial involving 160 patients, termed OPTIMA (OPtimal Target Identification via Modeling of Arrhythmogenesis), to demonstrate the effectiveness of CDTs in guiding atrial ablation procedures [5].
+Table 2: Summary of the related representative existing surveys. DL: deep learning.
+Source & Publish Year Scope Gap
+Corral et al. (2020) [31] cardiac digital twins challenges and clinical applications
+Coorey et al. (2022) [34] cardiac digital twins challenges and clinical applications
+Bear et al. (2015) [35] ECG forward problem not for inverse problem
+Bergquist et al. (2021) [36] ECG forward problem not for inverse problem
+Gulrajani et al. (1998) [37] ECG forward and inverse problem conventional solutions
 Uhlmann et al. (2014) [38] inverse problem not for ECG; conventional solutions
@@
 Genzel et al. (2022) [40] inverse problem not for ECG; mainly DL-based solutions
-Arridge et al. (2019) [44] inverse problem not for ECG; mainly DL-based solutions
-Sameni et al. (2010) [45] fetal ECG analysis fetal ECG; partially includes inverse problem
-Bear et al. (2015) [46] ECG forward problem not for inverse problem
-Hong et al. (2020) [47] DL-based ECG analysis not for inverse problem
-Liu et al. (2021) [48] DL-based ECG analysis not for inverse problem
+Siontis et al. (2021) [41] DL-based ECG analysis not for inverse problem
 Somani et al. (2021) [42] DL-based ECG analysis not for inverse problem
-Bifulco et al. (2021) [49] translational applications of computational modelling translational applications
-Cantwell et al. (2015) [50] ECG inverse problem mainly focusing on the inference of ATM and CV
-Cluitmans et al. (2015) [51] ECG inverse problem mainly focusing on conventional methods
-Cluitmans et al. (2018) [52] ECG inverse problem mainly focusing on the validation and clinical application
-Yadan et al. (2023) [53] ECG inverse problem mainly focusing on the identification of AF drivers
-Hernandez et al. (2023) [54] ECG inverse problem mainly focusing on the atria
+Bifulco et al. (2021) [43] ECG inverse problem applications of computational modeling
+Macleod et al. (1998) [6] ECG inverse problem conventional ECGI solutions
+Cluitmans et al. (2018) [7] ECG inverse problem conventional ECGI solutions and applications
+Hernandez et al. (2023) [44] ECG inverse problem atrial ECGI
+Loewe et al. (2022) [45] ECG inverse problem partially EP parameter estimation
+1.2 Study Inclusion and Literature Search
+In this work, we aim to provide readers with a survey of the state-of-the-art solutions to ECG inverse problems.
+There exist several modalities for cardiac electrical activity assessment, such as ECG, BSP, electrogram (EGM), electro-anatomic mapping (EAM), vectorcardiography (VCG).
+However, in the context of solving the ECG inverse problem, we usually only involve non-invasive ECG and BSP as the observed data.
+Note that compared to standard 12-lead ECG, BSPM is, however, not yet widely available as a diagnostic modality [46].
+For cardiac activity representation, three major source models are employed, i.e., surface potential models, activation models, and transmembrane potential (TMP).
+For model personalization, the EP parameters to be estimated can be conductivities, conduction velocities (CVs), earliest activation sites (EASs), tissue excitability, etc.
 To ensure comprehensive coverage, we have screened publications mainly from the last 24 years (2000-2024) related to this topic.
@@
 We followed their recent works to obtain state-of-the-art methodologies and clinical applications.
-Both peer-reviewed journal papers and conference papers were included here.
+Both peer-reviewed journal and conference papers were included here.
 In the way described above, we have collected a comprehensive library of more than 200 papers.
-Note that we picked the most detailed and representative ones for this review when we encountered several papers from the same authors about the same subject.
+Note that we normally only picked the most detailed and representative ones for this review when we encountered several papers from the same authors about the same subject.
 1.3 Existing Survey from Literature
@@
 One can see that the scopes of current review works are different from ours, though with partial overlaps.
-For example, there are several existing reviews focusing on summarizing solutions of inverse problems but not for ECG data [40, 41, 42, 44, 43].
-Alternatively, there exist few surveys on the ECG analysis, but not focusing on the inverse problem [46, 48, 1].
+For example, several surveys have been conducted to summarize the current challenges and existing or potential clinical applications of ECGI [7] or CDTs [31, 34].
+Furthermore, there are several existing reviews focusing on summarizing solutions of inverse problems but not for ECG data [38, 39, 40].
+Alternatively, there exist few surveys on the ECG analysis, but not focusing on the inverse problem [41, 42].
 Bifulco et al.
@@
 Although Macleod et al.
-[7] and Dossel et al.
-[39] performed a methodology survey on the ECG inverse problem, we focus on the developments proceeding their reviews.
-Cantwell et al.
-[50] mainly summarized the techniques for local activation time annotation and conduction velocity estimation in cardiac mapping.
-We provide a comprehensive and up to date review of ECG inverse inference, encompassing both its recent methodological aspects and potential clinical applications.
+[6], Cluitmans et al.
+[47] and Hernandez et al.
+[44] performed a methodology survey on the ECG inverse problem, they only summarized the ECG inverse inference studies for cardiac source reconstruction, i.e., ECGI.
+Loewe et al.
+[45] partially reviewed the ECG inverse inference works on EP parameter estimation in terms of cellular and tissue levels.
+Here, we provide a comprehensive and up-to-date review of ECG inverse inference in a broader range, encompassing both cardiac source reconstruction and EP parameter estimation.
+The communities engaged in these two areas of research do not completely overlap and sometimes even use different terminologies.
+For example, in the narrow definition of the inverse-ECG problem, i.e., ECGI, the forward model typically refers to the relationship between cardiac sources and ECG data.
+In contrast, within the model personalization community, the forward model often pertains to the forward EP model.
+This survey aims to bridge the gap between these communities, fostering conversation and collaboration.
 Figure 2: Summary of the ECG inverse inference methodologies. MCMC: Markov Chain Monte Carlo.
@@
 2.1 Definition of the Problem
-The ECG inverse problem refers to the mathematical task of determining the electrical activity within the heart based on measurements of the ECG signals recorded on the body surface.
-It involves applying a quasi-static approximation of Maxwell’s equations on the electromagnetic field between the heart and the torso to establish a relationship between the ECG data and the cardiac electrical excitation.
-Solving these equations numerically on subject-specific heart-torso models yields a linear measurement model,
+The ECG inverse problem involves inferring internal cardiac information from surface ECG measurements. This problem can be formulated in two ways:
+  1. 1.
+    Reconstruction of cardiac sources (ECGI): It aims to reconstruct the cardiac electrical activity from the recorded ECG or body surface potentials (BSPs) for a non-invasive visualization of cardiac activity.
+The relationship between the ECG data $\boldsymbol{y}(\boldsymbol{s},t)$ and the cardiac sources $\boldsymbol{u}(\boldsymbol{s},t)$ at position $\boldsymbol{s}$ and time instant $t$ is typically modeled as a linear system:
     $\boldsymbol{y}(\boldsymbol{s},t)=\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)+\epsilon,$ (1)
-where $\boldsymbol{y}(\boldsymbol{s},t)$ and $\boldsymbol{u}(\boldsymbol{s},t)$ refer to the vector of ECG signals and cardiac electrical parameters at position $\boldsymbol{s}$ and time instant $t$, $\boldsymbol{H}$ denotes the forward matrix, and $\epsilon$ represents the noise term that accounts for measurement and modelling errors.
-Based on this physical system modelled by PDEs and/or ordinary differential equations (ODEs), the typical solution of the problem is to inversely estimate $\boldsymbol{u}(\boldsymbol{s},t)$ that minimizes the fitting of the observational data $\boldsymbol{y}(\boldsymbol{s},t)$,
-$\min_{\boldsymbol{u}(\boldsymbol{s},t)}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2},$ (2)
-which however is ill-posed.
-The inverse problem can exhibit one of three characteristics: it may have multiple solutions (non-unique), no solution (non-existence), or solutions that are highly sensitive to small perturbations in the measured data (unstable) [55].
-Moreover, it is typically influenced by measured input data that is frequently corrupted by random or stochastic noise.
-To solve the ECG inverse problem, the state-of-the-art approaches can be coarsely separated into two kinds: deterministic and probabilistic methods [56], as presented in Fig.
-3.
+    where $\boldsymbol{H}$ is approximated as a forward matrix, and $\epsilon$ accounts for noise.
+The goal is to estimate $\boldsymbol{u}(\boldsymbol{s},t)$ by minimizing the fitting of the observational data $\boldsymbol{y}(\boldsymbol{s},t)$:
+    $\min_{\boldsymbol{u}(\boldsymbol{s},t)}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}.$ (2)
+    Note that this linear model specifically applies to potential-based ECG inverse problems.
+Activation-based models, which often involve non-linear relationships, are not directly addressed by this formulation [48, 49, 50].
+  2. 2.
+    Estimation of EP model parameters: It focuses on personalize cardiac models by estimating patient-specific EP parameters within a mathematical model of cardiac electrophysiology, such as the Aliev-Panfilov (AP) model [51] or Eikonal model [1].
+The AP model can be defined as,
+    $\frac{\partial u}{\partial t}$ $=\nabla\cdot(\boldsymbol{D}\nabla\boldsymbol{D})-cu(u-\theta)(u-1)-\frac{\partial u}{\partial z},$ (3)
+    $\frac{\partial z}{\partial t}$ $=\epsilon(u,z)\left(-z-cu(u-\theta-1)\right),$
+    where $u$ is the normalized transmembrane action potential, $z$ is the recovery current, $\boldsymbol{D}$ is the diffusion tensor, $c$ controls the repolarization, and $\theta$ is tissue excitability parameter.
+The Eikonal equation describes the propagation of the activation wavefront:
+    $\begin{cases}|\nabla\tau(\boldsymbol{s})|=\dfrac{1}{v(\boldsymbol{s})},&\boldsymbol{s}\in\Omega,\\ \tau(\boldsymbol{s})=\tau_{0}(\boldsymbol{s}),&\boldsymbol{s}\in\Gamma_{0},\end{cases}$ (4)
+    where $\tau(\boldsymbol{s})$ represents the activation time at point $\boldsymbol{s}$ within the cardiac tissue domain $\Omega$, while $v(\boldsymbol{s})$ denotes the local CV.
+The boundary conditions are defined by the set of EASs $\Gamma_{0}$, with $\tau_{0}(\boldsymbol{s})$ specifying the activation times at these sites.
+Taking the Eikonal model as an example, its parameter optimization objective function can be formulated as,
+    $\min_{\boldsymbol{v},\tau_{0}}\|\boldsymbol{y}_{\text{ECG}}-\boldsymbol{y}_{\text{model}}(\boldsymbol{v},\tau_{0})\|_{2}^{2}+\lambda\mathcal{R}(\boldsymbol{v}),$ (5)
+    where $\boldsymbol{y}_{\text{ECG}}$ is the measured signal, while $\boldsymbol{y}_{\text{model}}(\boldsymbol{v},\tau_{0})$ represents the simulated ECG signals using the forward EP model with CVs $\boldsymbol{v}$ and EASs $\tau_{0}$; $\mathcal{R}(\boldsymbol{v})$ imposes certain constraints on the estimated parameters, and $\lambda$ is a regularization term.
+Both are essential for creating CDTs, but they may involve different methods.
+The state-of-the-art approaches can be coarsely separated into two kinds: deterministic and probabilistic methods for both [52], as presented in Fig.
+2.
 Deterministic approaches in cardiac electrophysiology involve minimizing a cost function that quantifies the discrepancy between the observed data and the model predictions.
 Rather than aiming solely to minimize a cost function, probabilistic models embrace uncertainty and provide a probabilistic distribution over possible outcomes.
+We will review these two approaches separately and summarize their application to different formulations of ECG inverse inference.
+While we categorize the methodologies into deterministic and probabilistic, it is important to note that these categories often overlap, which will be emphasized when introducing these methods.
 2.2 Deterministic Estimation Approaches
 2.2.1 Spatial Regularization
-To stabilize and improve the solution of the ECG inverse problem, spatial regularization methods are employed [57].
+To stabilize and improve the solution of the ECG inverse problem, spatial regularization methods are widely employed for ECGI [53].
 These methods introduce spatial constraints to promote smoothness in the estimated electrical activities across the heart surface at different orders of derivatives [54].
@@
 For example, the Tikhonov regularization, including both zero-order and first-order variants, is the most widely used regularization technique [25, 59, 60, 61, 50, 62].
-It minimizes the mean squared error from the body-heart transformation while simultaneously penalizing the L2 norm of the inverse HSP solution.
+It minimizes the mean squared error from the body-heart transformation while simultaneously penalizing the L2 norm of the inverse heart surface potential (HSP) solution.
 The objective function can be formulated as,
@@
 where the notation $\|\cdot\|^{2}$ represents the L2 norm, $\lambda_{Tikh}$ denotes the regularization coefficient, and $\Gamma$ is the operator constraining the HSP solution $u(\mathbf{s},t)$, which is an identity matrix for zero-order regularization and a spatial gradient operator for first-order regularization to increase the spatial smoothness of the inverse solution.
-This dual process effectively suppresses the unreliable components in the solution and enhances overall smoothness for improved accuracy.
+This dual process effectively suppresses the unreliable components in the solution and enhances overall smoothness for improved accuracy [63].
+The choice of $\lambda_{\text{Tikh}}$ is crucial for enhancing reconstruction quality, and the L-curve method is commonly employed to determine the optimal value [64, 63].
 TSVD regularization method directly filters small singular values of the transfer matrix $\boldsymbol{H}$ and thus be robust to noises.
 TTLS regularization is effective in processing geometric errors that exist in the ECG inverse problem [56].
-L1-norm and TV regularization techniques are useful for addressing specific challenges in the inverse problem, such as noise suppression, sparsity, and edge preservation.
+L1-norm and TV regularization can tackle specific challenges in the inverse problems, such as noise suppression, sparsity, and edge preservation [65].
+Note that although these methods have been extensively applied to HSP based solutions, they can be applied to other linear inverse problem representation, such as TMPs [66].
 2.2.2 Spatiotemporal Regularization
-While spatial regularization methods are effective and widely used in ECG inverse problems, they often treat the problem as a static spatial problem and ignore the temporal dynamics of cardiac electrical activity.
-In some cases, this assumption may be reasonable, especially for certain types of cardiac abnormalities or when the temporal variations in the electrical activity are relatively slow compared to the sampling rate.
-However, in situations where the temporal dynamics are critical or when the electrical activity changes rapidly, temporal regularization methods that explicitly consider the time-varying nature of the problem may be more appropriate.
-These methods can help capture dynamic changes in electrical activity over time and provide more accurate and physiologically relevant solutions.
-To fuse the space-time correlations into the inverse ECG problem, spatiotemporal regularizations are employed for robust estimation of HSP.
-For instance, besides employing Tikhonov regularization in the spatial domain, Messnarz et al.
-[68] introduced additional temporal constraints by assuming that cardiac electric potentials are non-decreasing during the depolarization phase.
+While spatial regularization methods are effective and widely used in ECGI, they often overlook the temporal dynamics of cardiac electrical activity.
+To fuse the space-time correlations into the ECGI, spatiotemporal regularizations are therefore employed for robust estimation of cardiac sources.
+For instance, Twomey regularization was proposed as a modification of Tikhonov regularization to incorporates temporal a priori information [67, 68].
+It adjusts the second term in Eq.
+6 to minimize the difference between the solution and the prior estimate from the previous time instant.
+Alternative, one could introduce several additional constrains, including temporal constraints, to increase robustness to changes in the values of the regularization parameters [69, 70, 71].
 Yao et al.
-[69] proposed a spatiotemporal regularization method by regularizing both the spatial and temporal smoothness in the inverse ECG modelling, with the following objective function,
+[72] proposed a spatiotemporal regularization method by regularizing both the spatial and temporal smoothness, with the following objective function,
 $\min_{\boldsymbol{u}(s,t)}\sum_{t}\bigg{\{}\|\boldsymbol{y}(\boldsymbol{s},t)-\boldsymbol{H}\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}+\lambda_{s}^{2}\|\Gamma\boldsymbol{u}(\boldsymbol{s},t)\|_{2}^{2}$ (7)
@@
 where $\lambda_{s}$ and $\lambda_{t}$ are the spatial and temporal regularization terms, respectively, and $\omega$ denotes the selected time window to incorporate the temporal correlation.
-Splines-based methods can use spline functions to represent the electrical activity on the heart surface over time.
+Recently, frequency-guided model have be widely applied for imaging the spatiotemporal behavior of the heart [48, 73, 74, 49].
+Alternatively, splines-based methods can use spline functions to represent the electrical activity on the heart surface over time.
 They take advantage of the smoothness properties of splines to regularize the inverse solution, effectively capturing both spatial and temporal variations in the electrical potentials [75].
-A Kalman filter reformulation can incorporate temporal information to regularize the inverse problem [71, 72].
+Erem et al.
+[76, 77] employed the temporal spline model to reconstruct TMP and ATM and localize pacing sites from BSP.
+However, their approach faced the challenge of determining the optimal number of spline functions in advance.
+To overcome this limitation, more flexible adaptive splines have been proposed to reconstruct cardiac sources, enabling data-driven selection of spline functions from a large set of basis elements through a stepwise process [75, 78].
 Recently, to achieve regularization over space and time simultaneously, Cluitmans et al.
-[73] firstly applied regularization in the sparse wavelet domain for solving the ECG inverse problem, namely wavelet-promoted spatiotemporal regularization.
-Their experimental results indicate that the sparse representation of the cardiac source is beneficial for improving reconstruction accuracy in electrocardiographic imaging.
+[79] applied regularization in the sparse wavelet domain for reconstruct activation wavefronts and epicardial potential (EpiP), namely wavelet-promoted spatiotemporal regularization.
+Note that several studies have employed spatiotemporal Bayesian maximum a posteriori (MAP) or Kalman filters for ECGI [80, 81, 82], which however belong to probabilistic estimation and will be discussed further in Sec.
+2.3.1 and 2.3.2.
+Figure 3: Example of the physics-based model based on graph convolutional neural networks (CNNs) for the inverse inference of heart surface potential from body surface potential (BSP) [83].
+Illustrations designed referring to Jiang et al.
+[83] and Bear et al.
+[84].
 2.2.3 Model-based Regularization
@@
 In contrast, model-based regularization can integrate prior physiological knowledge about the electrical propagation inside the heart to constrain the solution space [85].
+Compared to the above regularization methods that primarily focus on ECGI, model-based methods have been widely applied to both ECGI and EP parameter estimation.
+For ECGI, model-based regularization can enhance the representation of cardiac electrical activity by incorporating detailed models of activation and wave propagation.
 For instance, step jump functions and logistic functions have been employed to model the activation of action potential [86, 87].
 Parameterized curves have been utilized to represent the wavefront velocity as trigonometric functions, while the potential has been modelled using the step response of a second-order linear system [68].
-The ECG inverse problems can also be formulated via level-sets by evolving a boundary from an initial region to minimize a filtered residual error, allowing for iterative refinement of the solution.
-The major advantage of level-set-based methods is that no isotropy assumptions are required, allowing for more flexible modelling and accurate representation of complex geometries and phenomena within the heart [77].
+The ECGI problems can also be formulated via level-sets by evolving a boundary from an initial region to minimize a filtered residual error, allowing for iterative refinement of the solution.
+The major advantage of level-set based methods is that no isotropy assumptions are required, allowing for more flexible modeling and accurate representation of complex geometries and phenomena within the heart [88, 89].
+Cluitmans et al.
+[85] introduced a electrophysiology based regularization to model the cardiac surface potential using a set of basis vectors.
+The epicardial potential is then reconstructed as a sparse combination of these vectors, based on EP principles.
 To estimate transmural potentials throughout the myocardium, 3D biophysical EP simulation models have been employed to offer spatiotemporal constraints of the inverse issue [90, 91, 92].
-While model-based regularization is effective in regularizing the ill-posed problem, their employed physiological models are developed for the general population [79, 74].
-Specifically, the parameters of these models, which are typically linked to tissue characteristics (excitability and contractility of the heart muscle), are often assigned values commonly used in the literature.
-The general physiological models could overly constrain subject-specific information, leading to potential model errors and model-data mismatch [81].
-For instance, the parameters controlling the configuration of the transmembrane potential (TMP) will differ throughout the heart, depending on whether the underlying tissue is in a healthy or diseased state.
-Therefore, Erem et al.
-[82] investigated how the solution changes when the assumption of uniform TMP amplitudes throughout the heart fails in the presence of infarction.
-Similarly, Xu et al.
-[81] and Dhamala et al.
-[83] emphasized the importance of uncertainty estimation in providing interpretable reconstructions of cardiac electrical activity.
+Similar strategies have been previously employed for predicting onset activation location [93] and reconstrucing 3D activation time from BSP [94, 48, 95].
+Furthermore, there exist a few model-based ECGI studies that optimize the model parameters via maximizing the correlation between simulated and measure BSP, which sometimes could be quite computational expensive [96, 97].
+For EP parameter estimation, model-based regularization serves a different purpose.
+It aims to identify patient-specific parameters rather than applying generalized physiological models that are commonly used for ECGI [92].
+This is generally achieved by optimization methods based on gradient descent [98, 99].
+Furthermore, Chinchapatnam et al.
+[100] proposed an adaptive zonal decomposition iterative algorithm to estimate the local CV and conductivity within the Eikonal model.
+Similarly, Sermesant et al.
+[101] defined the regularization term as the energy of a simplified dynamical electromechanical model of the heart to estimate EASs and local CV.
 2.2.4 Data-Driven Neural Networks
-The notable rise in computational resources and the vast increase in available data have facilitated the development of data-driven models.
-Conventional deterministic models for ECG inverse problems typically rely on physics-based mathematical models, such as the bidomain, the monodomain, or the Eikonal model, to describe the propagation of electrical activity through the heart.
-In contrast, data-driven deterministic models can employ neural networks to directly learn the mapping from ECG data to the underlying electrical sources without relying on explicit physics-based equations.
-These models are highly flexible and can capture complex, non-linear relationships in the data.
-They do not rely on strong assumptions about the underlying physiology, making them suitable for cases where the true underlying model is not well-known, varies across patients, or displays unexpectedly abnormal behavior.
-Therefore, they have been employed to solve the ECG inverse problem.
-For example, Meister et al.
-[84] designed a deep learning method based on graph convolutional neural networks (CNNs) to estimate the depolarization patterns in the myocardium with scars.
-By leveraging the inherent graph structure of the cardiac anatomy, graph CNNs enable the extraction of spatial dependencies and patterns, facilitating a more nuanced representation of electrocardiographic data.
+The notable increase in computational resources and available data have facilitated the development of data-driven models for the ECG inverse inference.
+For ECGI, traditional methods often rely on regularization techniques and/ or biophysical models to stabilize the solution, but these approaches may struggle with accuracy and robustness in the presence of data noise or model errors.
+In contrast, data-driven neural networks can learn complex mappings between BSP and cardiac source directly from data, without the need for explicit biophysical modeling [102].
+They do not rely on strong assumptions about the underlying physiology, making them suitable for cases where the true underlying model is not well-known, varies significantly across patients, or exhibits unexpected abnormal behavior.
+Therefore, they offer a promising alternative for solving the ECG inverse problem, particularly in scenarios with high variability or uncertainty in patient-specific cardiac physiology.
+The pioneering data-driven ECGI work was conducted by Yang et al.
+[103], who utilized the 12-lead ECG and convolutional neural networks (CNNs) to localize the origins of premature ventricular contractions (PVCs).
+They developed two distinct CNN models: one to classify the PVC origin across different segments and another to distinguish between endocardial (Endo) and epicardial (Epi) origins.
+Later, Meister et al.
+[104] designed a deep learning method based on graph CNNs to estimate the depolarization patterns in the myocardium with scars.
+Similary, Mu et al.
+[105] combined graph CNN and iterative soft threshold algorithm to reconstrcuct TMP and ATM from BSP.
+By leveraging the inherent graph structure of the cardiac anatomy, graph CNNs enable the extraction of spatial dependencies and patterns, facilitating a more nuanced representation of ECG data.
 Tenderini et al.
-[85] utilized a PDE-aware deep learning model for epicardial potential (EpiP) reconstruction from ECG data.
+[106] utilized a PDE-aware deep learning model for EpiP reconstruction from ECG data.
 Camara et al.
-[86] converted the BSP into an image represented by a 3D matrix and then employed CNNs to predict the locations of AF drivers from BSP.
-Neumann et al.
-[87] employed a reinforcement learning model to achieve multi-physics computational model personalization.
-Similarly, the time delay neural network can employ both current and previous timestep values of BSP to predict the HSP [88, 89, 90].
-However, these purely data-driven models generally lack interpretability and may underperform in scenarios with limited data due to their data-intensive training requirements.
-Figure 4: Example of the physic-informed neural networks for the reconstruction of activation time map.
-Here, the Eikonal equation has been employed as an additional physical constraint [91].
-Image modified from Sahli [91] with permission.
-Figure 5: Example of the physics-based model based on graph convolutional neural networks (CNNs) for the inverse inference of epicardial potential from body surface potential (BSP) [92].
-Illustrations designed referring to Jiang et al.
-[92] and Bear et al.
-[93].
+[107] converted the BSP into an image represented by a 3D matrix and then employed CNNs to predict the locations of atrial fibrillation drivers from BSP.
+Cheng et al.
+[108] employed fast iterative shrinkage/thresholding network to reconstruct TMP from ECG.
+Chen et al.
+[109] tested the feasibility of several different neural networks, including CNN, fully connected neural network, and long short-term memory, for the reconstruction of EpiP and ATM as well as the localization of pacing sites.
+Furthermore, time delay neural networks have been employed to extract both current and previous timestep values of BSP to predict the HSP [110, 111, 112].
+Interestingly, Karoui et al.
+[111] found that direct data-driven ATM reconstruction outperformed the two-stage approach, which first reconstructs HSP via a neural network and then estimates the ATM from the HSP using intrinsic deflection time.
+This is because small fluctuations in the reconstructed HSP can lead to errors in the subsequent computation of ATM.
+For EP parameter estimation, the first deep learning based EP parameter personalization model was proposed by Neumann et al.
+[113], which however belongs to probabilistic estimation approaches (see Sec.
+2.3).
+This is because they reformulated the EP parameter estimation problem into a Markov decision process and applied reinforcement learning to personalize the global CVs and Purkinje network [114].
+Later, Giffard et al.
+[115] employed transfer learning to estimate local CVs and pacing sites from BSP for personalized cardiac resynchronization therapy (CRT).
+Nevertheless, these purely data-driven models generally lack interpretability and may underperform in scenarios with limited data due to their data-intensive training requirements.
 2.2.5 Physics-Informed Neural Networks
@@
 This can be achieved either via introducing underlying physical constraints as an additional loss [117, 118], or designing specialized network architectures that satisfy the physics laws [83, 119].
-For instance, Xie et al.
-[95] combined the data-driven loss and physics-based loss to inversely predict the heart-surface potentials from BSP and evaluated their framework in a 3D torso-heart geometry.
-Ye et al.
-[96] proposed a spatial-temporally adaptive PINN framework for solving both ECG forward and inverse problems over 3D bi-ventricular models.
-Sahli et al.
-[98] used PINNs in forward mode to estimate activation time maps (ATMs) and conduction velocity (CV) maps, as shown in Fig.
-4.
+For ECGI, Xie et al.
+[117] combined the data-driven loss and physics-based loss to inversely predict the HSP from BSP and evaluated their framework in a 3D torso-heart geometry.
+Kashtanova et al.
+[120] proposed a APHYN-EP model, which decomposing the transmembrane dynamics into a physical term and a data-driven term, respectively.
+The data-driven deep learning component is tailored to capture information that the incomplete physical model fails to represent.
+This concept was later formally defined as the hybrid neural EP model by Jiang et al.
+[121].
+Instead of introducing additional loss, Jiang et al.
+[83] proposed a spatial-temporal graph CNN-based model, where the geometry-dependent physics between BSP and HSP can be explicitly modelled via a bipartite graph over their graphical embeddings, as presented in Fig.
+3.
+For EP parameter estimation, Ye et al.
+[118] proposed a spatial-temporally adaptive PINN framework for both ECG simulation and AP model parameter estimation.
+Herrero et al.
+[119] developed a PINN model for action potential reconstruction and EP parameter estimation from sparse amounts of EP data.
+Instead, Sahli et al.
+[122] used PINNs in the forward model to interpolate activation time maps (ATMs) and estimate CV maps from a few sparse activation measurements.
 They incorporated a physics-informed regularization loss prescribed by the Eikonal equation, which describes the relationship between ATMs and the spatial gradient of CV.
-Similarly, Herrero et al.
-[97] developed a PINN model for action potential reconstruction and EP parameter estimation from sparse amounts of EP data.
-Although they did not inversely infer ATMs from BSP or ECG, a similar idea can be employed to solve the ECG inverse inference problem.
-Instead of introducing additional loss, Jiang et al.
-[92] proposed a spatial-temporal graph CNN-based non-Euclidean encoding-decoding model, where the geometry-dependent physics between BSP and HSP can be explicitly modelled via a bipartite graph over their graphical embeddings, as presented in Fig.
-5.
-In general, the use of PINN in cardiac EP is mainly limited to 1D/ 2D in silico data, as it could be computationally expensive in realistic 2D/ 3D settings [97, 96].
+Although they did not inversely infer ATMs from BSP or ECG, a similar idea can be employed to achieve the ECG inverse inference.
+In general, PINNs for ECG inverse inference are mainly used for 1D/2D in silico data, as they face significant challenges, such as hyperparameter tuning, managing complex geometries, and spectral bias.
+These issues become more pronounced in realistic 2D/3D settings, due to their complex geometries and need for accurate spatial derivatives, leading to high computational requirements [119, 118].
 2.3 Probabilistic Estimation Approaches
-Unlike deterministic regularization techniques, probabilistic or stochastic estimation methods provide solutions in the form of probability distributions.
-Consequently, multiple potential solutions can be derived from the distributions, effectively addressing the inherent uncertainties within the data and the models relevant to the ECG inverse inference [19].
-Specifically, the inverse problem can be formulated as the estimation of a Bayesian posterior probability distribution function (pdf), which can be defined as,
-$p(\boldsymbol{u}|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{u},\epsilon)p(\boldsymbol{u}),$ (5)
-where $p(\boldsymbol{u})$ is the prior, and the likelihood $p(\boldsymbol{y}|\boldsymbol{u},\epsilon)$ can be modeled as a normal distribution $\mathcal{N}(\boldsymbol{H}\boldsymbol{u},\epsilon\boldsymbol{I})$ [99].
-One can predict the mode of the posterior distribution via maximum a posteriori (MAP) estimator, which is a point estimator that finds the parameter value that maximizes the posterior pdf given the observed data.
-When dealing with the complex dynamics of cardiac electrical activity, the Kalman filter serves as a specialized instance of the MAP estimator, enabling the integration of spatio-temporal information for the ECG inverse inference [100].
-Alternatively, to estimate the entire posterior distribution, one can employ Markov Chain Monte Carlo sampling or variational inference for approximating analytically intractable posterior distributions.
-These approaches allow for a more comprehensive understanding of the cardiac electrical activity, taking into account the inherent variability and errors in the data and models, ultimately leading to more reliable and interpretable results.
+Unlike deterministic techniques, probabilistic estimation methods provide solutions in the form of probability distributions.
+Specifically, these approaches can naturally incorporate spatiotemporal information by using spatiotemporal priors or modeling state transitions [123].
+In fact, deterministic methods such as Tikhonov regularization can be formulated as a special and simplified case of Bayesian MAP estimation under linear models with Gaussian priors [123].
+In these cases, the regularization parameter effectively represents prior knowledge about the model.
+Probabilistic methods extend this by not only incorporating such prior information but also by providing a distribution of possible solutions.
+This distribution allows for comprehensive statistical analysis and helps address uncertainties inherent in the data and models relevant to ECG inverse inference [19].
+2.3.1 Bayesian MAP Estimation
+In Bayesian MAP estimation, the solution is selected by maximizing the posterior probability of the sources given the measurements.
+For ECGI, this maximization aims to find the most likely cardiac source distribution that matches both prior physiological knowledge and observed ECG measurements, given a specific forward model [124].
+Greensite et al.
+[125] proposed using spatial and spatiotemporal Bayesian MAP methods for bioelectric (including ECG) inverse problems.
+Similarly, Zhou et al.
+[126] used a spatiotemporal Bayesian MAP approach to reconstruct EpiP and localize activation origins from BSP, which was later extended to the endocardial surface [127].
+Instead of relying solely on BSP, Serinagaoglu et al.
+[128] utilized additional statistical prior information about the EpiP distribution, along with sparse EpiP measurements obtained from multielectrode coronary venous catheters, to enhance EpiP reconstruction within a Bayesian framework.
+Erenler et al.
+[123] assumed a time-invariant model when employing the Bayesian MAP for EGM reconstruction, which performed worse than Kalman filter solutions that consider the spatiotemporal variations of EGM.
 2.3.2 Kalman Filter
-The Kalman filter (KF) is a well-established tool for state estimation in dynamic systems.
-In the context of cardiac modelling, it implicitly estimates the posterior distribution of the underlying electrical activity given the observed ECG measurements.
-This is achieved by modelling the cardiac system as a high-dimensional stochastic state-space model, where the KF or its extensions can be effectively applied [79].
-The key concept involves estimating the state $\boldsymbol{u}(\boldsymbol{s},t)$ by sequentially updating its estimate in the presence of noisy observations.
-The state transition is described as follows,
-$\boldsymbol{u}(\boldsymbol{s},t)=\boldsymbol{F}\boldsymbol{u}(\boldsymbol{s},t-1)+\boldsymbol{w}_{t},$ (6)
-where $\boldsymbol{F}$ refers to the state transition matrix (STM) and $\boldsymbol{w}_{t}$ is the process noise vector which models uncertainties and errors arising from the STM [101].
-Then, the KF calculates the Kalman gain and uses it to update the state estimate based on observed data, considering spatio-temporal variations within cardiac source distributions.
-It provides not only a point estimate of the state but also a covariance matrix that represents the uncertainty associated with the estimate.
-By modelling both the dynamics of the system and the measurement noise probabilistically, KF inherently accounts for uncertainty in the estimation process.
+The Kalman filter (KF) is a well-established method for estimating states in dynamic systems and has been used to improve ECGI by incorporating temporal information [129, 68, 123].
+It implicitly estimates the electrical activity of the heart based on noisy ECG measurements by modeling the heart as a state-space system [92, 82].
+The KF updates estimates of the state $\boldsymbol{u}(\boldsymbol{s},t)$ using the state transition matrix $\boldsymbol{F}$ and process noise $\boldsymbol{w}_{t}$ [130],
+$\boldsymbol{u}(\boldsymbol{s},t)=\boldsymbol{F}\boldsymbol{u}(\boldsymbol{s},t-1)+\boldsymbol{w}_{t}.$ (8)
+The filter calculates a point estimate and a covariance matrix representing the uncertainty.
+It effectively handles uncertainty by modeling both the system dynamics and measurement noise.
 To reduce computational complexity, researchers often focus on temporal relationships among spatially well-correlated regions, such as neighbouring regions or leads belonging to the same activation wavefront.
@@
 Other formulations, like the Duncan and Horn formulation of the KF used by Berrier et al.
-[71], focus on reconstructing specific electrograms.
+[129], focus on reconstructing specific EGMs.
 Wang et al.
 [92], in an attempt to integrate general prior knowledge with subject-specific data, used a statistical perspective to explicitly account for both model and data errors.
-They represented the cardiac electrophysiological system in a state-space form and conducted inverse inference of 3D TMP distributions via the unscented KF.
+They represented the cardiac EP system in a state-space form and conducted inverse inference of 3D TMP distributions via the unscented KF.
 A similar strategy was employed to define both the electrical and mechanical measurements as the state variables, which can be estimated via the unscented KF [131].
-While the KF is a powerful tool for state estimation in cardiac modelling, it involves large matrices operation and may suffer from the initialization problem.
-2.3.2 Markov Chain Monte Carlo Sampling
-Figure 6: Example of the MCMM application on the inverse inference of cardiac activation properties from ECG.
+Huang et al.
+[132] proposed a data-driven Kalman filter model, consisting of a state transfer network and a Kalman gain network.
+Their method can consider the dynamic activation process of the cardiomyocyters when recovering TMP from BSP.
+While the KF is a powerful tool for state estimation in cardiac modeling, it involves large matrices operation and may suffer from the initialization problem.
+For EP parameter estimation, to the best of our knowledge, only one study has used a reduced-order unscented KF to estimate cardiac tissue conductivity for personalizing a cardiac EP model [133].
+In constrast, KF has been widely employed for the cardiac electromechanical model parameter estimation, such as contractility [134, 135].
+Figure 4: Example of the MCMM application on the inverse inference of earliest activation sites (EASs) and CV from ECG.
 Here, MCMC has been used for the sampling of parameter sets of the population [1].
@@
 [1] with permission.
-Current probabilistic inference approaches are generally based on Markov Chain Monte Carlo (MCMC) sampling methods to handle complex probabilistic models and uncertainty associated with ECG data [104].
+2.3.3 Markov Chain Monte Carlo Sampling
+Probabilistic estimation of EP parameters is crucial for model personalization and uncertainty quantification.
+Many probabilistic inference methods use Markov Chain Monte Carlo (MCMC) sampling to handle complex models and uncertainties associated with ECG data [136].
 For instance, Camps et al.
-[103] presented a sequential Monte Carlo approximate Bayesian computation-based model for the inference of ventricular activation properties.
+[1] presented a sequential Monte Carlo approximate Bayesian computation-based model for the inference of ventricular activation properties, as shown in Fig.
+4.
 Rahimi et al.
 [137] applied a hierarchical Bayesian inference to estimate the statistical parameters in the posterior pdf, which can be calculated via MCMM sampling methods.
-However, direct MCMC sampling of the pdf of the parameters can be impractical, as it involves repeated evaluations of the posterior pdf which consists of computationally expensive simulation model.
+However, direct MCMC sampling of the pdf of the parameters can be impractical, as it involves repeated evaluations of the posterior pdf involving computationally expensive simulation processes.
 To accelerate sampling, various hybrid sampling techniques have been developed using information about the target pdf such as gradient and Hessian matrix [138], which however are challenging to extract when dealing with complex simulation models.
@@
 Instead of completely replacing the sampling of the exact posterior pdf, Dhamala et al.
-[83] proposed a two-stage model to integrate Gaussian process surrogate modelling of the posterior pdf in accelerating sampling.
-Also, one could directly develop a surrogate model for the posterior pdf of simulation model parameters, eliminating the necessity for additional MCMC sampling of the computationally intensive original pdf [104].
+[141] proposed a two-stage model to integrate Gaussian process surrogate modeling of the posterior pdf in accelerating sampling.
+Also, one could directly develop a surrogate model for the posterior pdf of simulation model parameters, eliminating the necessity for additional MCMC sampling of the original computationally-intensive pdf [136].
 2.3.4 Variational Inference
-Variational inference (VI) approximates the posterior distribution with a simpler variational distribution by solving an optimization problem.
-This optimization process is often computationally more efficient than MCMC sampling as it does not require repeated sampling from the posterior distribution.
-Therefore, the application of VI has been widely adopted in solving the ECG inverse problem [74].
-For instance, Xu et al.
+For ECGI, unlike conventional regularization techniques, variational inference (VI) provides a probabilistic framework that can quantify uncertainty in the reconstructed potentials.
+Specifically, it can incorporate prior knowledge about the spatial and temporal dynamics of cardiac activity, improving the robustness of the solution against noisy data.
+For example, Ghimire et al.
+[142] introduced a Bayesian framework to jointly infer the posterior distribution of TMP and the errors in the prior EP model from ECG data under the constraint of a 3D EP simulation model.
+Xu et al.
 [143] introduced a Bayesian inference framework for robust transmural electrophysiological imaging.
-Ghimire et al.
-[74] proposed a Bayesian model for the estimation of the prior model error while reconstructing TMP under the constraint of a 3D EP simulation model.
-These conventional VI methods are generally optimized using techniques such as gradient descent and expectation-maximization (EM) algorithms.
-Recently, deep learning-based probabilistic methods have served as an emerging alternative to conventional methods for modelling complex dynamics of cardiac electrical activity [109, 110, 111].
+Recently, deep learning-based probabilistic methods have served as an emerging alternative to conventional methods for modeling complex dynamics of cardiac electrical activity [144, 145].
 They can leverage deep neural networks to approximate the posterior distribution of the model parameters or latent variables, providing faster and more accurate approximations.
-For example, Ghimire et al.
-[109] proposed a deep generative model to reconstruct cardiac transmembrane potential from ECG data.
-Bacoyannis et al.
-[112] employed a variational auto-encoder (VAE) model to reconstruct the activation pattern of the myocardium with various local wall thicknesses as thin walls indicated infarct areas.
+Specifically, variational auto-encoder (VAE), as a extension of traditional VI, is quite useful to handle complex, high-dimensional data during the inverse inference.
+For example, Bacoyannis et al.
+[102] employed a VAE model to reconstruct the activation pattern of the myocardium with various local wall thicknesses as thin walls indicated infarct areas.
 Bacoyannis et al.
 [146] employed conditional VAE to generate activation maps from combined BSP and conductivity maps.
-Li et al.
-[114] designed a VAE-based deep computational model for the inverse inference of ventricular activation properties, i.e., root node (RN) position and CVs.
-Dhamala et al.
-[115] employed a graph convolutional VAE to allow generative modelling of non-Euclidean data for cardiac model personalization.
-Dhamala et al.
-[110] proposed a generative VAE for parameter estimation of a personalized cardiac model.
-Yang et al.
+For EP parameter estimation, VI approximates the posterior distribution of EP parameters with a simpler variational distribution by solving an optimization problem.
+Similar to ECGI, VAE has also been widely used for EP parameter estimation, such as the inverse inference of ventricular activation properties, i.e., EASs and CVs [147].
+Graph convolutional VAE can further enhance this estimation by effectively handling non-Euclidean data structures, i.e., cardiac meshes, allowing for more accurate modeling of spatial relationships within the heart [148].
+Specifically, Dhamala et al.
+employed graph convolutional VAE to embed the high-dimensional optimization into low-dimensional latence space for the estimation of tissue excitability [149].
+Recently, Yang et al.
 [145] proposed Bayesian PINN to solve linear or nonlinear PDEs with noisy data for both forward and inverse problems.
 They employed the Hamiltonian Monte Carlo or the variational inference for the estimation of the posterior distributions.
-Figure 7: Example of the conditional-VAE model for the inverse inference of myocardial infarction (MI) [116]. Image adapted from Li et al. [116] with permission.
+This approach could be effectively adapted to tackle the ECG inverse problem in the future.
 2.3.5 Surrogate Model for Bayesian Optimization
-To accelerate the inverse inference of parameters, surrogate models, especially statistical emulation with Gaussian processes (GP), have been employed in cardiovascular fluid dynamics [117], the pulmonary circulatory system [118], ventricular mechanics [119], and LV cardiac dynamics [117].
+To accelerate the inverse inference of parameters, surrogate models, especially statistical emulation with Gaussian processes (GP), have been employed in cardiovascular fluid dynamics [150], the pulmonary circulatory system [151], ventricular mechanics [152], and LV dynamics [150].
 It involves approximating the computationally intensive mathematical model (the simulator) with a computationally efficient statistical surrogate model (the emulator) [150].
-This is achieved through a combination of extensive parallelization and GP regression, allowing for optimal utilization of computational resources prior to the patient arrival at the clinic.
-Here, GP regression is a non-parametric Bayesian approach used for regression tasks where the model learns from the observed data to make predictions.
+This could be achieved through a combination of extensive parallelization and GP regression, allowing for optimal utilization of computational resources prior to the patient arrival at the clinic.
+GP regression is a non-parametric Bayesian approach used for regression tasks where the model learns from the observed data to make predictions.
 It does not assume a specific parametric form for the underlying function but instead models the function as a distribution over functions, which is determined by the observed data points.
@@
 Recently, the GP surrogate model started to be used in the ECG inverse inference by Pezzuto et al.
-[120], where the GP surrogate model with Bayesian optimization has been used for the inverse inference of RNs.
-These strategies offer a mid-way solution between pure sampling methods using mechanistic models and machine learning-based solutions as the variational inference.
-They retain the flexibility and interpretability of pure sampling methods while replacing the mechanistic model with a faster emulator, which can be machine learning-based.
-3 Uncertainty Quantification for the ECG Inverse Problem
-Solving the ECG inverse problem relies on many subject-specific measurements and modelling assumptions, which introduce various sources of uncertainty in the predictions.
-Uncertainty quantification aims to analyze how models respond to inherent variability in input parameters.
-ECG inverse problems are often ill-posed and lack uniqueness in their solutions.
-However, clinicians typically require a single prediction for practical decision-making purposes in patient care scenarios.
-Therefore, it is essential to quantify and mitigate the effects of different types of uncertainties to ensure the robustness and clinical applicability of ECG inverse solutions.
-3.1 Measurement Acquisition Uncertainty
-ECG inverse inference requires subject-specific EP measurements, which can be easily distorted by noise artifacts, such as motion artifacts.
-EP data typically exhibit sparsity, with measured values associated with a higher degree of uncertainty [121].
-For example, noise, a ubiquitous challenge in real-world measurements, can significantly distort the recorded ECG signals, introducing uncertainty into the source localization.
-To reduce the effects of measurement noise on the inverse problem, Aydin et al.
-[101] used a Kalman filter-based approach.
-Dogrusoz et al.
-[122] employed Bayesian estimation for solving the inverse ECG problem by applying various prefiltering methods to reduce the effects of noise.
-Similarly, Neumann et al.
-[123] utilized stochastic parameter estimation, aggregating probabilities estimated under various noise levels to derive a robust parameter estimate without explicitly determining the level of noise in the measurement.
-Moreover, there is ambiguity in matching the measurement points in the BSPM recording with the computational mesh derived from imaging data due to the non-synchronized BSPM-imaging acquisition [107].
-The heart position is also susceptible to various sources of movement artifacts, including respiration and cardiac contraction, introducing geometric errors into the inverse problem of ECG solutions [124].
-Therefore, the geometric error in the anatomical modelling (forward model) should also be considered, as it will in turn cause estimation errors in the ECG inverse inference.
-3.2 Anatomical Modeling Uncertainty
-The anatomical modelling is normally started with image segmentation, and then numerical solution techniques will be used to solve the forward ECG problem [58].
-Therefore, the precision of the anatomical model employed, hinging on segmentation accuracy [125, 126], inter-subject anatomical variations [127], structural remodelling of the myocardium [116], and mesh resolution [128], profoundly influences the reliability of inverse solutions.
-Incorporating a prior model can mitigate uncertainties, and a combination of prior models may be preferred for estimating complex source structures [105].
-Also, one can disentangle the inter-subject anatomical variation to mitigate its effect on the ECG inverse problem [127].
-Furthermore, biventricular pacing and scar size, often encountered in clinical scenarios, demand meticulous modelling for a faithful representation [129].
-Besides the heart, the torso geometry accuracy can also subtly influence the inverse solution, as the EP measurement is generally performed on the body surface [130].
-In most studies, each torso surface node often corresponds to an electrode, and thus insufficient electrodes may lead to non-representative mesh nodes, requiring interpolation techniques for solving the inverse problem [131].
-Dogrusoz et al.
-[131] investigated the effects of interpolation on the inverse problem of ECG.
-There also exist many studies exploring the effects of cardiac position inside the torso on the inverse inference [132, 133].
-To address uncertainties, advanced regularization techniques that encompass prior knowledge, noise reduction strategies, and advanced data assimilation methods show promise [58, 134].
-As the field advances, careful consideration of these effects becomes paramount for robust and clinically meaningful results in solving the ECG inverse problem.
-Figure 8: The cardiac digital twin generation workflow combining cardiac imaging and ECG data, with ECG inverse inference for personalization. Image modified from Li et al. [116] with permission.
-3.3 Computational Modeling Assumption Uncertainties
-Model assumptions are necessary simplifications or idealizations of reality that are made to facilitate mathematical or computational modelling.
-These assumptions may include assumptions about tissue properties, boundary conditions, geometry, and physiological processes.
-In ECG inverse inference, assumptions are often made about the homogeneity of tissue properties, the behavior of electrical signals within the heart, and the relationship between measured ECG signals and underlying electrical activity [126, 135, 136].
-While model assumptions are essential for simplifying complex systems, they can also introduce uncertainties and errors into the modelling process.
-For example, assuming uniform tissue properties throughout the heart and torso may lead to inaccuracies in the predicted electrical activity, especially in regions with significant heterogeneity [130, 53].
-Similarly, assuming specific boundary conditions or physiological behaviors may not fully capture the complexity of real-world scenarios, leading to discrepancies between model predictions and observed data.
-ECG inverse inference also depends heavily on the assumptions on the initial distribution of heart electrical potential [135].
-Determining the initial potential distribution at the onset of the cardiac cycle is particularly challenging due to the difficulty in accurately pinpointing the anatomical site of the earliest excitation in the heart, often associated with the intricate Purkinje networks [95].
-Therefore, sophisticated modelling strategies are required to adequately account for the influence of these networks on ECG inverse inference [137, 138].
-Once uncertainties are quantified, strategies can be employed to mitigate their impact on model predictions.
-The process of balancing model assumptions and uncertainty is often iterative, requiring continuous refinement and validation of models against experimental or clinical data.
-This iterative refinement process helps researchers to build more robust and reliable models over time.
+[153], where the GP surrogate model with Bayesian optimization has been used for the estimation of EASs.
+Earlier work by Konukoglu et al.
+[139] demonstrated Bayesian estimation using polynomial chaos for EASs and conductivity estimation.
+These strategies offer a mid-way solution between pure sampling methods using EP models and machine learning-based solutions as the variational inference.
+They retain the flexibility and interpretability of pure sampling methods while replacing the model with a faster emulator, which can be machine learning-based.
 3 Dataset and Evaluation
+Figure 5: Illustration of Utah Torso Tank, with an isolated canine heart. Image adapted from Bergquist et al. [36] with permission.
+Figure 6: Illustration of in-vivo evaluation based on the rabbit undergoing the ventricular pacing from various locations. Illustrations designed referring to Han et al. [95].
 3.1 Dataset
-The solution to the ECG inverse inference problem is typically implemented within a cardiac digital twin platform, as shown in Fig.
-8.
-Therefore, for evaluation purposes, access to both anatomical data from cardiac imaging and electrical activity recording data from ECG or BSP mapping is necessary.
-While there are many public cardiac imaging data [139, 140, 141, 142, 143, 144] and ECG recordings [145, 146, 147, 148], the evaluation of ECG inverse inference methods requires paired imaging and ECG/ BSP mapping data from the same patient, which is less publicly available.
-UK Biobank (UKB) [149] and China Kadoorie Biobank (CKB) [150, 151] provide paired imaging and ECG dataset.
+To evaluate the solution of ECG inverse problem, both anatomical data from cardiac imaging and electrical activity recording data from ECG or BSP are necessary, as shown in Fig.
+1.
+While there are many public cardiac imaging data [154, 155, 156, 157] and ECG recordings [158, 159, 160], the evaluation of ECG inverse inference methods requires paired imaging and signal data from the same patient, which is less publicly available.
+UK Biobank [161] and China Kadoorie Biobank [162, 163] provide paired imaging and ECG dataset.
 MyoFit46 includes 500 participants of approximately 75 years+ to undertake high-density surface ECGI and cardiac MRI [164].
@@
 Experimental Data and Geometric Analysis Repository (EDGAR) is another public dataset for the application and validation of ECGI techniques [166].
-It includes paired BSPM and endocardial/ epicardial recordings from both animals (dog, pig, etc) and human and torso tank datasets.
-Additionally, information regarding the cardiac position within the torso is essential, often requiring torso image data.
-Several studies have attempted to reconstruct the torso directly from standard cardiac imaging for the localization of electrodes, which was challenging due to partial torso shape information [155, 156].
-There are also many validation studies conducted in ex vivo torso tanks [157, 158, 159], in vivo large animal models [160, 161, 93], or in patients [162, 25].
+It includes paired BSPM and endocardial/ epicardial recordings from both animals and torso tank datasets.
+Accurate cardiac positioning within the torso is also crucial and often requires torso imaging data for precise localization.
+Several studies have attempted to reconstruct the torso directly from standard cardiac imaging for the localization of electrodes, which was challenging due to partial torso shape information [167, 168, 169].
+There are also several validation studies conducted in ex vivo torso tanks [36, 170, 171], in vivo large animal models (rabbit, canine, swine, etc) [172, 173, 84, 95, 174], or in human [175, 26].
 Fig.
-9 and Fig.
-10 provide the two most commonly used ex vivo torso tanks.
-Figure 9: Illustration of Utah Torso Tank, with an isolated canine heart. Image adapted from Bergquist et al. [157] with permission.
-Figure 10: Illustration of Bordeaux Torso Tank, with a pig heart perfused in Langendorff mode with a 108-electrode sock placed over the ventricles.
-Image adapted from Bear et al.
-[158] with permission.
-Figure 11: Illustration of different cardiac source models in ECG.
+5 provides the one of the most commonly used ex-vivo torso tanks.
+Fig.
+6 presents an in-vivo rabbit model for the evaluation of non-invasive 3D activation sequence reconstruction.
 3.2 Evaluation
-The validation of the inverse problem involves a diverse set of cardiac source models, forward EP model formulations, and inverse methods, as presented in Table 3.
-For simplicity, we organized the validation according to the estimated cardiac source models, including surface potential models, activation/ recovery models, and transmembrane voltage models [52].
-Fig.
-11 provides three representative examples corresponding to the three source models.
-Except for a few studies, all the others employed BSP as the input representation/ measurement for cardiac electrical activity assessment.
-One can see that most studies employed computer models for in silico evaluation [66, 163] and/ or performed evaluation on the basis of torso-tank experiments with isolated animal hearts, such as rabbit [164], canine [77, 73, 92], and swine [165, 166, 167].
-There are also few studies evaluated using in-vivo human subjects, which however usually employed non-simultaneous invasive recordings, such as EGMs [168, 169] or bipolar voltages maps [81, 74].
+Table 3: Summary of selected ECG inverse inference works for ECGI.
+EGM: electrograms; OVA: origin of ventricular activation; VT: ventricular tachycardia; MI: myocardial infarction; AF: atrial fibrillation; PVC: premature ventricular contraction; EpiP/ EndoP: epicardial/ endocardial potential.
+Source & Publish year Data Signal # Electrodes Target Inverse inference method
+Deterministic estimation Shou et al. (2008) [56] in silico BSP N/A EpiP truncated total least squares
+Wang et al. (2009) [59] in silico BSP N/A EpiP Tikhonov regularization
+Wang et al. (2013) [176] in silico BSP 400 Myo ischemia, TMP Tikhonov regularization, total-variation
+Cluitmans et al. (2018) [79] 3 canines BSP 192 ATM, EpiP Wavelet regularization
+Zhou et al. (2018) [65] 7 VT patients ECG; BSP 120 OVA multi-linear regression; L1-norm regularization
+Kalinin et al. (2019) [177] in silico BSP N/A EpiP, EndoP Tikhonov regularization
+Kara et al. (2019) [60] in silico BSP 32/64/128/512/ 1024 Myo ischemia, EpiP Tikhonov regularization
+Bear et al. (2020) [61] 4 canines/swines BSP 192/128 EpiP, ATM Tikhonov regularization
+Rababah et al. (2021) [62] 6 canines/swines BSP 192/128 EpiP, ATM Tikhonov regularization
+Schuler et al. (2021) [50] in silico BSP 200 EpiP, TMP, ATM Tikhonov regularization
+Messnarz et al. (2004) [70] in silico BSP 62 TMP spatiotemporal regularization
+Erem et al. (2013) [76] 3 subjects BSP 120 EndoP, ATM, pacing sites temporal spline model
+Erem et al. (2014) [77] in silico; 1 subject BSP 65 TMP, ATM temporal spline model
+Zhou et al. (2016) [73] 7 AF patients ECG 208 atrial EndoP spatiotemporal frequency based model
+Yao et al. (2017) [178] 4 MI patients BSP N/A MI, EGM spatiotemporal regularization
+Coll et al. (2017) [74] in silico; 1 canine BSP 300, 354 T-wave alternans, EpiP, EndoP spatiotemporal spectral and frequency based model
+Yang et al. (2018) [49] in silico; 4 swines BSP 208, 128 ATM spatial gradient sparse in frequency domain
+Li et al. (2001) [93] in silico BSP 200 OVA model optimization based method
+Tilg et al. (2002) [179] 2 patient ECG 55, 49 ATM model optimization based method
+Calderero et al. (2005) [88] 1 canines BSP 711 activation wavefront level set
+Liu et al. (2006) [48] in silico BSP 64/96/128/155/ 200 ATM biophysical model-based method
+Han et al. (2008) [95] 4 rabbits BSP 53 ATM biophysical model-based method
+Chavez et al. (2015) [89] 3 canines BSP N/A Myo ischemia iterative level set
+Cluitmans et al. (2017) [85] 3 canines BSP 184-216 EpiP physiology-based regularization
+Yang et al. (2017) [103] in silico; 4 PVC patients ECG 3 origin of PVC data-driven neural network
+Karoui et al. (2019) [110] in silico BSP N/A EpiP spatial adaptative time delay neural network
+Alawad et al. (2019) [180] in silico; 3 patients ECG N/A OVA transfer learning
+Gyawali et al. (2019) [181] 39 VT patients ECG N/A OVA sequential autoencoder
+Cheng et al. (2021) [108] in silico BSP N/A TMP, MI shrinkage-thresholding network
+Mu et al. (2021) [105] in silico BSP N/A TMP, MI, ATM shrinkage-thresholding + graph CNN
+Camara (2021) et al. [107] in silico BSP 64 AF driver data-driven neural network
+Karoui et al. (2021) [111] in silico BSP N/A atrial ATM data-driven neural network
+Meister et al. (2021) [104] 16 swines ECG N/A ATM graph convolutional regression
+Chen et al. (2022) [109] 5 swines BSP 239 EpiP, ATM, pacing sites data-driven neural network
+Tenderini et al. (2022) [106] in silico ECG 9 EpiP PDE-aware deep learning model
+Jiang et al. (2024) [121] in silico; 3 patients BSP N/A EpiP, EndoP neural state-space modeling
+Lian et al. (2024) [182] in silico; 4 patients ECG 9 MI frequency-enhanced network
+Jiang et al. (2022) [83] 1 canine BSP 192 EpiP physics embedded graph CNN
+Xie et al. (2023) [117] in silico BSP N/A EpiP physics-constrained network
+Probabilistic estimation Serinagaoglu et al. (2006) [128] 1 canine BSP 771 EpiP Bayesian MAP
+Wang et al. (2010) [183] in silico BSP N/A TMP, MI Bayesian MAP
+Zhou et al. (2018) [127] in silico; 3 patients BSP 120 EndoP, pacing sites sparse Bayesian MAP
+Zhou et al. (2018) [126] in silico; 3 VT patients BSP 120 EpiP spatio-temporal Bayesian MAP
+Wang et al. (2009) [92] in silico; 1 MI patient BSP 700/330, 123 TMP physiological-constrained Kalman model
+Huang et al. (2022) [132] in silico BSP N/A TMP data-driven Kalman filtering network
+Wang et al. (2010) [184] 1 canine BSP N/A EpiP variational-form-based regularizers
+Wang et al. (2011) [185] 1 canine BSP N/A EpiP variational regularizers
+Xu et al. (2014) [186] in silico; 6 patients BSP N/A MI variational Bayesian
+Figuera et al. (2016) [53] in silico BSP N/A atrial EpiP Bayesian MAP, etc
+Ghimire et al. (2019) [142] in silico; 2 patients ECG N/A TMP Bayesian model
+Ozkoc et al. (2021) [187] 1 canine BSP 192 EpiP Bayesian MAP
+Li et al. (2024) [188] in silico ECG 10 MI multi-modal VAE
+Table 4: Summary of the selected ECG inverse inference works for EP parameter estimation. HF: heart failure; APD: action potential duration; MS: Mitchell-Schaeffer; AP: Aliev-Panfilov.
+Source & Publish year Data Signal Target EP model Inverse inference method
+Deterministic estimation Chinchapatnam et al. (2008) [100] in silico; 1 patient N/A local CV; conductivity Eikonal adaptive zonal iterative model
+Sermesant et al. (2012) [101] 2 patients ECG local CV, EAS Eikonal model-based regularization
+Potse et al. (2014) [189] 2 HF patients ECG EAS, conductivity, etc Ten Tusscher and Panfilov model-based regularization
+Kahlmann et al. (2017) [190] in silico ECG Purkinje network Eikonal parallel simplex optimization
+Sanchez et al. (2018) [191] 6 patients ECG EAS, conductivity, etc Ten Tusscher and Panfilov iterative method
+Lee et al. (2019) [192] in silico ECG global CV reaction-Eikonal rule based model
+Pezzuto et al. (2021) [193] 11 HF patients ECG global CV, EAS Eikonal propagation model
+Grandits et al. (2021) [98] in silico; 1 patient ECG EAS Eikonal topological gradient
+Costa et al. (2022) [8] 6 pigs ECG global CV, APD, conductivity reaction-Eikonal total activation time fit
+Grandits et al. (2023) [99] 1 patient ECG local CV, EAS, ATM Eikonal Geodesic backpropagation
+Neumann et al. (2016) [114] in silico ECG global CV, Purkinje network Eikonal reinforcement learning
+Giffard et al. (2018) [115] 20 patients BSP local CV, pacing site, ATM MS transfer learning
+Herrero et al. (2022) [119] in silico; in vitro ECG APD, excitability and diffusion coefficients AP PINN
+Ye et al. (2023) [118] in silico ECG tissue excitability AP PINN
+Probabilistic estimation Zettinig et al. (2014) [46] in silico ECG electrical diffusivity MS statistical polynomial regression
+Gillette et al. (2021) [2] 1 patient ECG global CV, Purkinje network, APD, conductivity reaction-Eikonal Saltelli sampling
+Camps et al. (2021) [1] in silico ECG global CV, EAS Eikonal Monte Carlo Bayesian inf
+Camps et al. (2024) [194] in silico ECG global CV, Purkinje network Eikonal Monte Carlo Bayesian inf
+Zaman et al. (2021) [136] in silico; 3 VT patients ECG tissue excitability AP Bayesian active learning
+Dhamala et al. (2020) [149] in silico; 3 VT patients ECG tissue excitability AP generative VAE
+Li et al. (2022) [147] 1 patient ECG global CV, EAS Eikonal conditional VAE
+Konukoglu et al. (2011) [139] in silico; 1 patient BSP EAS, conductivity Eikonal polynomial chaos based Bayesian inf
+Dhamala et al. (2017) [195] in silico; 2 VT patients ECG tissue excitability AP spatially adaptive surrogate model
+Pezzuto et al. (2022) [153] in silico ECG EAS Eikonal GP based Bayesian optimization
+The validation of the inverse problem involves a diverse set of targeted cardiac sources or EP parameters, input signal, forward EP model formulations, and inverse inference methods, as presented in Table 3 and Table 4.
+Note that the tables only provide a selection of representative studies and are not an exhaustive list of all available studies.
+One can see that BSP has been widely used for reconstructing cardiac electrical activity, while ECG is commonly used for EP parameter estimation.
+Several studies have shown that the number of electrodes greatly affects the accuracy of cardiac source reconstruction [48, 60].
+Instead of using fixed number of electrodes, Cluitmans et al.
+[85] proposed determining the number based on the torso size of the individual.
+Alternatively, Ondrusova et al.
+[196] proposed a two-step single dipole based inverse inference model, initially using all electrodes and then refining with a subset selected by significance via a greedy algorithm.
+For EP parameter estimation, the AP and (reaction-)Eikonal models are frequently used due to their balance between computational efficiency and accurate representation of cardiac activation dynamics.
+The Mitchell-Schaeffer (MS) model, while less commonly used, offers simplicity and is suitable for specific scenarios requiring less complexity [46, 115].
+Most studies employed computer models for in silico evaluation [50, 177] and/ or performed evaluation on the basis of torso-tank experiments with isolated animal hearts, such as rabbit [197, 95], canine [88, 79, 83], and swine [104, 198, 49, 109].
+Also, there exist several in-situ animal studies, where the potential of the heart and body surfaces were simultaneously recorded in animals, such as Auckland pig or Utah/ Maastricht dog data [84, 128, 173].
+There are also few studies evaluated using in-vivo human subjects, which however usually employed non-simultaneous invasive recordings, such as EGMs [193] or bipolar voltages maps [186, 142].
 The comparison with the published invasive measurements in animals or humans has also been utilized as an alternative evaluation [25].
-Table 3: Summary of the ECG inference works.
-EP: electrophysiology; AS: activation sequence; EGM: electrograms; HF: heart failure; HS: healthy subject; EpiP/ EndoP: epicardial/ endocardial potential; RN: root node; VCG: vectorcardiographic; AP: Aliev-Panfilov; LSTM: long short-term memory; FCN: fully connected neural network; TV: total-variation.
-Source & Publish year Data Input Target EP model Inverse inference method
-Ramanathan et al. (2004) [24] 1 HS; 3 patients BSP EpiP, EGM, isochrones boundary element Tikhonov regularization + generalized minimal residual algorithm
-Shou et al. (2008) [60] in silico BSP EpiP finite element truncated total least squares
-Wang et al. (2009) [63] in silico BSP EpiP finite element Tikhonov regularization
-Jiang et al. (2009) [170] in silico BSP EpiP boundary element subspace preconditioned LSQR
-Sarikaya et al. (2010) [59] 1 canine BSP EpiP N/A genetic algorithm with conventional regularization
-Wang et al. (2010) [171] 1 canine BSP EpiP finite element variational-form-based regularizers
-Wang et al. (2011) [172] 1 canine BSP EpiP finite element variational regularizers
-Cluitmans et al. (2013) [168] 1 patient BSP EpiP N/A realistic data and VCG based optimization
-Zemzemi et al. (2013) [173] in silico BSP EpiP bidomain Steklov-Poincare formulation + gradient descent method
-Zemzemi et al. (2014) [174] in silico BSP EpiP bidomain iterative Kozlov-Maz’ya-Fomin
-Erem et al. (2014) [82] in silico; 1 HS BSP EpiP, EndoP N/A temporal spline model
-Yao et al. (2017) [175] 4 MI patients BSP EGM, MI N/A spatiotemporal regularization
-Cluitmans et al. (2017) [176] 3 canine BSP EpiP AP physiology-based regularization
-Cluitmans et al. (2018) [73] 3 canine BSP EpiP, ATM, beat origin N/A multitask elastic-net-based method
-Kalinin et al. (2019) [163] in silico BSP EpiP, EndoP bidomain single layer approach
-Karoui et al. (2019) [88] in silico BSP EpiP monodomain spatial adaptation of time delay neural network
-Bear et al. (2020) [65] 4 canine/swine BSP EpiP, ATM boundary element Tikhonov regularization
-Rababah et al. (2021) [67] 6 canine/swine BSP EpiP, ATM boundary element Tikhonov regularization
-Ozkoc et al. (2021) [177] 1 canine BSP EpiP boundary element Bayesian MAP
-Chen et al. (2022) [167] 5 swines ECG EpiP, ATM, RN N/A FCN, CNN, LSTM model
-Tenderini et al. (2022) [85] in silico ECG EpiP AP PDE-aware deep learning model
-Jiang et al. (2022) [92] 1 canine BSP EpiP AP physics embedded graph CNN
-Xie et al. (2023) [178] in silico BSP EpiP AP physics-constrained network
-Jiang et al. (2024) [179] in silico; 3 patients BSP EpiP, EndoP reaction-Eikonal model neural state-space modeling
-Messnarz et al. (2004) [68] in silico BSP TMP bidomain spatiotemporal regularization
-Wang et al. (2009) [79] in silico BSP TMP bidomain physiological-model-constrained statistical approach
-Wang et al. (2011) [180] in silico BSP TMP, MI bidomain physiologically-based constraints
-Potyagaylo et al. (2012) [181] in silico BSP TMP N/A iterative Gauss-Newton optimization
-Jiang et al. (2013) [182] in silico BSP TMP bidomain Support vector regression
-Lopez et al. (2014) [183] in silico BSP TMP monodomain genetic algorithm with Tikhonov regularization
-Xu et al. (2014) [81] in silico; 6 patients BSP MI N/A variational Bayesian with TV prior
-Zettinig et al. (2014) [38] in silico ECG Myo electrical diffusivity LB polynomial regression
-Ghimire et al. (2019) [74] in silico; 2 patients ECG TMP finite element Bayesian model
-Kara et al. (2019) [64] in silico BSP TMP, MI bidomain Tikhonov regularization
-Cheng et al. (2021) [184] in silico BSP TMP, MI bidomain shrinkage-thresholding network
-Mu et al. (2021) [185] in silico BSP TMP N/A shrinkage-thresholding + graph CNN
-Zhao et al. (2022) [186] in silico BSP TMP bidomain data-driven Kalman filtering network
-Li et al. (2024) [116] in silico ECG MI Eikonal multi-modal VAE
-Lian et al. (2024) [187] in silico; 4 patients ECG MI bidomain frequency-enhanced model
-Zhang et al. (2005) [164] 4 rabbits BSP ATM N/A N/A
-Calderero et al. (2005) [77] 1 canine BSP activation wavefront potential-based level sets
-Alawad et al. (2019) [188] in silico; 3 patients ECG pacing sites N/A L1-norm regularization
-Schuler et al. (2021) [66] in silico BSP ATM bidomain Tikhonov regularization
-Pezzuto et al. (2021) [169] 11 HF patients ECG ATM, RN Eikonal propagation model
-Meister et al. (2021) [165] 16 swine ECG ATM N/A graph convolutional regression
-Camps et al. (2021) [103] in silico ECG CV, RN N/A Monte Carlo Bayesian inf
-Li et al. (2022) [189] in silico ECG CV, RN Eikonal conditional VAE
-Camps et al. (2024) [138] in silico ECG CV, Purkinje network Eikonal Monte Carlo Bayesian inf
+As for inference target, for ECGI, there are three major cardiac source models, i.e., surface potential models (EpiP and EndoP), activation models, and transmembrane voltage models, as shown in Fig.
+7.
+Additionally, some studies have focused on reconstructing regions of electrical conduction abnormalities [178, 105, 188, 103] and the origins of ventricular activation [65, 180, 181].
+In general, reconstruction efforts have primarily focused on the ventricles, with relatively few studies addressing atrial imaging, as specifically highlighted in Table 3.
+This might be due to the atrial thinner walls, irregular anatomy, and more complex electrical activity [199].
+For EP parameter estimation, EAS (or Purkinje network), CV, conductivity, and tissue excitability have been mainly considered.
+Figure 7: Illustration of different cardiac source models in ECGI.
 3.2.1 Surface Potential Model
-The prevalence of potential map estimation on the cardiac surface can be attributed to its practicality and clinical significance.
-It provides a powerful and accessible means of understanding cardiac electrical behavior.
-Electrophysiological information reconstructed on the cardiac surfaces, such as the epicardium or endocardium, is relatively easier to obtain and is invaluable for clinical insights.
-Most surface potential inference work primary focused on the reconstruction of epicardial (outer) potential rather than endocardial (inner) potential.
-This is reasonable, as epicardial potential is more accessible for measurement compared to endocardial potential.
-As for the forward EP model, boundary element, finite element, and Aliev-Panfilov (AP) model have been widely used.
-4.2.2 Activation/ Recovery Model
-Activation and recovery time maps offer valuable static isochrone images depicting the sequence of cardiac electrical depolarization and repolarization.
-These isochrones can be applied in the clinic in various aspects: (a) identifying sites where activation initiates, as seen in focal arrhythmias, e.g., premature ventricular contractions (PVC) origin; (b) mapping reentry circuits, e.g., in macroreentrant atrial flutter; (c) detecting abnormalities in propagation, including areas with delayed activation [190] or those exhibiting slow or abnormal conduction, e.g., scar-related ventricular tachycardia (VT) substrate; and (d) pinpointing areas with recovery abnormalities, e.g., large recovery gradients [191, 192].
-To date, the majority of validation studies have concentrated on the accuracy of activation mapping, with only a few studies assessing recovery [76, 193].
-Besides, one could estimate activation parameters for personalization.
-The estimation of EP parameters from ECG or BSP is relatively rare, as these parameters are typically more detailed and are best captured through invasive electrophysiological studies or high-resolution mapping techniques.
-For example, Grandits et al.
-[194] employed Huber regularization to estimate simultaneously distributed fibre architecture and CVs from sparse contact recordings.
+As summarized in Table 3, surface potentials, specifically EpiP and EndoP, are the most common targets for cardiac source reconstruction.
+The prevalence can be attributed to its practicality and clinical significance, with EpiP being particularly favored as it is more accessible for measurement than EndoP.
+Furthermore, surface potentials usually require less computational power compared to full-volume models.
+They are generally faster to functionalize since they deal with surface data, and often rely on established methods and datasets.
+Nonetheless, the reconstruction of surface potentials can suffer from non-uniqueness, particularly if the input data is not sufficiently detailed.
+3.2.2 Activation Time Model
+ATM is commonly used to describe the propagation of electrical activity across the heart.
+It is a simplified model focused on the activation phase and assumes a smooth, single wavefront, which limits its applicability to more complex cardiac events, such as reentry and arrhythmias.
+The functionalization of ATM is relatively quick as it focuses on timing rather than detailed voltage dynamics, simplifying the modeling process.
+Most studies have reconstructed the ATM on the atrial and/or ventricular surfaces from BSP or ECG [179, 111, 79, 62].
+Instead, some studies have reconstructed the 3D cardiac activation sequence throughout the myocardial volume, evaluated both in silico [48, 50] and in vivo, including rabbit and canine models [95, 200, 174].
+For example, Schuler et al.
+[50] conducted an in-silico study to identify best configurations for reducing line-of-block artifacts in ATM reconstructed from EpiP or TMP.
+Han et al.
+[95] performed an in vivo validation study by comparing reconstructed activation imaging with 3D intracardiac mapping in a group of four healthy rabbits undergoing ventricular pacing from different locations, as presented in Fig.
+6.
+The reconstruction of ATM can be more unique in terms of identifying activation sequences but might lack the details needed to reconstruct the full electrical activity accurately.
 3.2.3 Transmembrane Potential Model
-Using the transmembrane potential (TMP) as a cardiac source model offers a highly accurate approximation of the cardiac actual electrical source [52].
-Estimating volumetric myocardial TMP dynamics is clinically significant as it unveils intricate arrhythmic details and intramural arrhythmogenic substrates.
-However, this necessitates complex modelling and interpretation of electrical activity within the 3D cardiac structure.
-Reconstructions of TMP support the identification of depolarized and repolarized regions in the heart and their temporal evolution.
-They also delineate areas of reduced amplitude or completely absent TMP, such as ischemic, fibrotic, or border-zone areas.
-Nevertheless, TMPs cannot be directly measured in the clinic, necessitating indirect validation via extracellular signals generated by TMP distributions.
-Surface potentials on the endocardial and epicardial surfaces, as well as activation and recovery times, can be calculated from TMP distributions and then measured in experiments.
+Using TMP as a cardiac source model provides an accurate approximation of the cardiac electrical activity, revealing critical arrhythmic details and intramural arrhythmogenic substrates [7].
+However, this approach is computationally demanding due to the need for complex 3D modeling and extensive data (volumetric mesh, fibre orientation, myocardial conductivities, etc), leading to longer setup and functionalization times.
+Despite these challenges, TMP reconstructions are valuable for identifying depolarized and repolarized regions and detecting ischemic, fibrotic, or border-zone areas.
+Since TMPs cannot be directly measured clinically, validation relies on extracellular signals generated by TMP distributions.
+Surface potentials and ATM can be derived from TMP distributions and then measured in experiments.
+3.2.4 EP Parameter
+As summarized in Table 4, the parameters to be personalized are directly tied to the forward EP models.
+For instance, the Eikonal model typically involves the personalization of parameters such as the EAS, CV, and tissue conductivity [100, 1, 194, 2].
+For example, Camps et al.
+[1] used dynamic time warping (DTW) and Pearson’s correlation coefficients to align simulated ECG with clinical acquired ECG, enabling the iterative updating of EASs and CVs, as shown in Fig.
+4.
+In contrast, the primary parameter optimized in AP models is usually tissue excitability [195, 149, 136, 118].
+However, these are not the only parameters that may require optimization.
+Depending on the specific application and model, additional parameters such as ionic channel properties, diffusion coefficients, or APD might also need to be personalized to accurately capture patient-specific cardiac behavior [2, 119].
+The choice of parameters to optimize is ultimately guided by the intended clinical application and the complexity of the model, balancing between computational feasibility and the need for precision in capturing the physiological phenomena of interest.
+3.3 Result
+Validating the results of ECG inverse inference is highly challenging due to the lack of a unified or standardized pipeline for ECGI or EP parameter estimation.
+The most commonly used evaluation metric for potential and activation models is the correlation coefficient (CC) [84, 63].
+For assessing the performance of pacing site, PVC origin, or cardiac excitation origin localization, the Euclidean or geodesic distance between the exact and reconstructed locations is typically employed [111, 109, 201, 202].
+As mentioned earlier, validation methods, source models, and the complexity of the forward model (e.g., homogeneous vs.
+inhomogeneous torso models) can vary significantly, leading to significant differences in results reported across studies.
+Generally, most previous studies have reported a median CC of approximately 0.7 for EpiP reconstruction, and this value can reach 0.8 to 0.9 for ATM and TMP reconstructions [109, 203].
+Pacing site localization errors can range from 2.0 to 20.0 mm, with a median distance of around 10 mm [109, 105, 202].
+A recent study compared the pacing site localization performance of dipole and potential-based source models using the public EDGAR dataset, reporting a median distance of 25.2 mm and 13.9 mm, respectively [202].
+For PVC origin localization, they reported errors of 30.2-33.0 mm and 28.9-39.2 mm for the dipole and potential-based model, respectively.
+The latest results for cardiac excitation origin localization reported an error of about 23.9 mm on clinical data after fine-tuning, and 44.0 mm without fine-tuning [121].
+This finding highlights the challenges of generalization when training ECGI reconstruction networks supervised by simulation data.
+Similar conclusions were drawn by Pilia et al.
+[201], who reported an error of about 1.5 mm on simulated data, which increased to 32.6 mm on unseen clinical data.
+ECGI has also been utilized to detect infarcted regions, achieving a average Dice score of 0.4 on simulated data [142, 188].
+EP parameter estimation involves determining intrinsic cardiac properties that are less directly observable, making its evaluation more complex.
+Furthermore, it requires sophisticated models to capture the underlying biophysical processes.
+Given these challenges, its validation often requires indirect methods, such as comparing simulation results against observed clinical measurement [1].
+Some studies use a combination of in-silico modeling and in vivo data, employing metrics such as root mean squared error (RMSE) and Dice coefficient for evaluation [136].
+These findings highlight the significant challenges in ECG inverse inference evaluation, emphasizing the need for standardized validation methods and improved model robustness.
 4 Clinical Application
-Inverse modelling of electrocardiographic signals can help to create a personalized virtual heart, which can be potentially applied for: (a) preoperative planning for cardiac surgery, (b) risk stratification and long-term monitoring, and (c) personalized disease insight and diagnosis.
-Figure 12: Virtual heart model guided preoperative planning for the personzalized catheter ablation of atrial fibrillation (AF) patients.
+While ECGI and EP parameter estimation focus on different aspects, i.e., detailed electrical imaging and personalized physiological modeling, respectively, they can be effectively integrated within the broader context of CDTs.
+CDTs leverage personalized data to create a comprehensive virtual heart model, encompassing detailed anatomical, electrophysiological, and dynamic information.
+It can not only optimize diagnostic and treatment planning, as with conventional ECGI, but also support long-term risk assessment and personalized disease management.
+Next, we will introduce the clinical applications of ECGI and the additional advantages brought by CDTs driven by personalized EP parameter estimation.
+For a more comprehensive overview of clinical applications, we refer readers to existing reviews on ECGI [7, 199] and CDTs [3, 204].
+4.1 Preoperative Planning for Cardiac Surgery
+The ability of ECGI to non-invasively characterize cardiac activation patterns is crucial for preoperative planning, such as catheter ablation.
+For example, ECGI has been effectively used to guide ablation in patients with persistent AF, achieving success rates of up to 85% of patients being free from AF 12 months post-procedure [205].
+However, ablation procedures are not totally free of risks, require significant resources, and are not suitable for all AF patients [206].
+Supporting the potential role of ECGI in patient selection, Meo et al.
+[207] demonstrated that ECGI could help identify patients likely to benefit from ablation.
+Similarly, Rodrigo et al.
+[208] showed that ECGI could stratify patients by projecting the complexity of electrical patterns onto the atrial surface, recommending ablation for those with less complex substrates.
+Despite these advances, clinical adoption of ECGI for AF ablation remains limited due to the incomplete understanding of AF mechanisms and ablation effects.
+In contrast, ECGI has gained more clinical acceptance for VT ablation [209].
+Netherless, in either the atrial or ventricle EP procedure setting, the cardiac imaging used in ECGI need to be better integrated with EP to optimize its usage.
+Figure 8: Virtual heart model guided preoperative planning for the personzalized catheter ablation of AF patients.
 Here, CARTO is the clinical electroanatomic navigation system.
@@
 [5].
+CDTs expand upon the capabilities of ECGI by providing a more comprehensive simulation environment for patient-specific planning.
+It enable the simulation of various surgical interventions, allowing surgeons to predict outcomes, optimize strategies, and assess risks in a fully controlled, risk-free setting.
+For instance, it has been applied to identify optimal infarct-related ablation targets for VT [210, 211].
+This can be accomplished by performing virtual multi-site delivery of electrical stimuli (pacing) from various bi-ventricular locations.
+Each location attempts to induce VT from a site positioned differently relative to the infarct area.
+The optimal ablation lesions can then be determined in each personalized virtual heart, rendering it non-inducible for VT from any pacing location.
+Therefore, it can terminate not only clinically manifested or induced VTs during the procedure but also all potential VTs arising from the post-infarction substrate.
+This includes VTs that may emerge after the initial ablation, potentially eliminating the need for repeated procedures and providing long-term freedom from VT for patients.
+Similarly, this technology allows for the preoperative identification of optimal ablation targets for AF [5], as shown in Fig.
+8.
+Additionally, CDTs hold the potential for the virtual implantation of cardiac devices, such as pacemakers, defibrillators, or transcatheter valves, to evaluate their placement and function [3].
+Complex surgical procedures, including coronary artery bypass grafting and valve repair or replacement, can also be simulated to optimize strategies, minimize risks, and improve patient outcomes [212, 213].
 Figure 9: Comparison of conventional clinical metric and virtual heart model based identification of patients at high sudden cardiac death (SCD) risk.
@@
 [4].
-5.1 Preoperative Planning for Cardiac Surgery
-Virtual heart models enable surgeons to simulate different surgical interventions and assess their feasibility and potential outcomes in a risk-free environment.
-For instance, utilizing personalized virtual-heart technology enables the identification of optimal infarct-related ablation targets for VT [193, 195, 196].
-This can be accomplished by performing virtual multi-site delivery of electrical stimuli (pacing) from various bi-ventricular locations.
-Each location attempts to induce VT from a site positioned differently relative to the infarct area.
-The optimal ablation lesions can then be determined in each personalized virtual heart, rendering it non-inducible for VT from any pacing location.
-Therefore, it can terminate not only clinically manifested or induced VTs during the procedure but also all potential VTs arising from the post-infarction substrate.
-This includes VTs that may emerge after the initial ablation, potentially eliminating the need for repeated procedures and providing long-term freedom from VT for patients.
-Similarly, this technology allows for the preoperative identification of optimal ablation targets for AF, as shown in Fig.
-12.
-It has proved invaluable in guiding ablation strategies, enabling clinicians to target specific areas responsible for atrial arrhythmias more effectively [3].
-For MI management, ECG inverse inference aids in delineating the extent and location of ischemic regions within the myocardium [74, 116].
-It can also be utilized to detect pulmonary vein (PV) drivers and thus assist in identifying patients with high acute success rates to PV isolation [197].
-Furthermore, surgeons can virtually implant cardiac devices such as pacemakers, defibrillators, or transcatheter valves to evaluate their placement and function.
-Additionally, complex surgical procedures such as coronary artery bypass grafting and valve repair or replacement can be simulated to optimize surgical strategies, minimize operative risks, and improve patient outcomes.
 4.2 Risk Stratification and Long-Term Monitoring
-In risk stratification, ECG inverse inference techniques contribute by unravelling intricate electrophysiological patterns that signify an increased risk of adverse cardiac events.
-For instance, Arevalo et al.
-[2] created a virtual-heart arrhythmia risk predictor (VARP) for sudden cardiac death (SCD) based on a virtual heart model, as shown in Fig.
-13.
-The heart model integrates the ventricular geometry of the patient, structural remodelling resulting from post-MI and electrical functions from sub-cellular to organ levels.
-Within each heart model, a virtual multi-site delivery of electrical stimuli originating from ventricular locations at varying distances to the remodelled tissue can be performed.
-This comprehensive evaluation allows for a thorough assessment of the patient’s heart propensity to develop ventricular arrhythmias related to MI.
-The VARP has been demonstrated to significantly outperform clinical metrics in predicting future arrhythmic events.
-The resilient and non-intrusive VARP approach holds promise in potentially averting SCD and steering clear of unnecessary implantable cardioverter-defibrillator (ICD) placements in patients with post-MI.
-This is particularly valuable in monitoring patients with known cardiovascular conditions, where early signs of deterioration can be identified, enabling timely intervention and personalized adjustments to the treatment plan [198, 199].
-This can be achieved by combining multiple sources (fitness trackers, wearables, and ’companion’ apps for drugs, etc.) to maintain and update cardiac digital twins and their underlying models.
-The integration of ECG inverse inference thus emerges as a powerful tool in risk assessment and continuous monitoring, contributing to more proactive and personalized cardiovascular care.
+By imaging and quantifying substrates associated with increased risk of adverse cardiac events, ECGI has been used for risk stratification [214].
+Early studies have identified key substrates for arrhythmias, such as slow discontinuous conduction pathways [215], abnormal repolarization [216], and steep repolarization gradients [217].
+CDT model provides a more comprehensive, personalized approach for risk stratification.
+It integrates the personalized geometry of the patient, structural remodeling resulting from diseases and electrical functions from sub-cellular to organ levels.
+For instance, a virtual-heart arrhythmia risk predictor (VARP) has been created for sudden cardiac death (SCD) based on a virtual heart model [4], as shown in Fig.
+9.
+Within the model, a virtual multi-site delivery of electrical stimuli originating from ventricular locations at varying distances to the remodelled tissue can be performed.
+It has been demonstrated to significantly outperform clinical metrics in predicting future arrhythmic events.
+The virtual heart model offers significant advantages over ECGI by allowing for non-invasive, personalized predictions of SCD risk, potentially preventing unnecessary implantable cardioverter-defibrillator (ICD) placements in post myocardial infarction patients.
+Moreover, it enables continuous monitoring of patients with known cardiovascular conditions, allowing for early detection of deterioration and timely adjustments to treatment plans [218, 219].
+By integrating data from multiple sources, such as fitness trackers, wearables, and companion apps, CDTs enhance the utility of ECG inverse inference, enabling more proactive and individualized cardiovascular care.
 4.3 Personalized Disease Insight and Diagnosis
-Numerous studies within the pathological validation section have delved into disease mechanisms using ECG inverse inference, proposing clinically relevant parameters [200, 201, 202].
-One of the earliest extensively clinically explored applications of ECG inverse inference is in identifying the specific mechanisms underlying arrhythmia onset and/ or perpetuation for ablation therapy [52].
-For instance, it was utilized to delineate the distinct activation patterns during arrhythmia (e.g., focal vs.
-re-entrant) [203, 204].
-Additionally, it has also been employed to pinpoint the drivers of this activation [205], which have demonstrated correlation with successfully ablated regions in patients with positive outcomes and are now the focus of direct targeting in clinical validation studies [206].
-Besides atria, ECG inverse inference has also been applied to identify and understand arrhythmogenic substrates for the onset and maintenance of VT [201, 207] and ventricular arrhythmias [208].
-It has showcased larger ventricular electrical dyssynchrony in CRT responders, improving patient selection over standard 12-lead ECG [209, 210].
-ECG inverse inference techniques are especially valuable in scenarios where invasive recordings are ethically constrained, such as with infants and pregnant women, offering insights into cardiac evolution.
-They can also be used for in silico clinical trials (see Section 6.3), which can particularly facilitate safe methods to explore treatment effects in patients with rare diseases and thus may allow for insights not possible in the current clinical trial practice [35].
-Certain diagnostic medical devices, like CardioInsight (Medtronic, USA), based on personalized models, have transitioned into industrial application and clinical usage [35].
-Within the CardioInsight mapping system, the cardiac electrical activity on its surface can be reconstructed from body surface potentials employing a customized model of the heart and torso of the target patient [211].
-While these retrospective studies contribute to understanding disease mechanisms and defining potential therapeutic markers, further exploration through extensive prospective studies is crucial to establish the clinical value of these insights for individual patients and their potential role in guiding diagnosis or patient selection.
+Numerous studies have demonstrated the potential of ECGI for understanding disease mechanisms and identifying clinically relevant parameters [220, 221, 222].
+One of the earliest and most explored clinical applications of ECGI is in the precise identification of arrhythmia mechanisms to guide ablation therapy [7].
+Also, ECGI has been used to delineate the distinct activation patterns during arrhythmia (e.g., focal vs.
+re-entrant) [223, 224], and to identify the primary drivers of these patterns, which often correlate with successful ablation targets [225, 73].
+Beyond atrial applications, it has also been effective in detecting and characterizing arrhythmogenic substrates that contribute to the onset and maintenance of VT [221, 226] and other ventricular arrhythmias [227] It has been shown to improve patient selection for CRT by identifying greater ventricular electrical dyssynchrony among responders compared to standard 12-lead ECG methods [228, 229].
+The growing commercial availability of ECGI systems, such as CardioInsight (Medtronic, USA), Corify Care, Vivo (Catheter Precision), and EP Solutions SA, further highlights its potential for clinical use [31, 75].
+At the same time, CDTs are transforming the landscape of in silico clinical trials, particularly for evaluating treatment effects in patients with rare diseases, where traditional trials have significant limitations [31].
+By enabling highly personalized simulations that capture complex cardiac dynamics and patient-specific characteristics, CDTs enhance the generalizability and reliability of trial results while reducing time and costs [194, 230].
+This technique accelerates drug development and provides deeper insights into drug effects, interactions, and mechanisms.
+While these studies have significantly advanced our understanding of cardiac disease mechanisms and identified potential therapeutic targets, further prospective studies are needed to validate these insights in clinical settings and explore their potential for guiding diagnosis and patient selection.
 5 Discussion and Future Perspectives
-6.1 Surrogate Model for Efficient ECG Simulation
+Here, we provide a structured discussion on the current challenges in ECG inverse inference and propose potential solutions to address these challenges.
+By evaluating uncertainties, exploring computationally efficient models, and integrating multimodal data, we aim to provide a comprehensive roadmap for future research in this field.
+5.1 Quantifying and Mitigating Uncertainties in ECG Inverse Inference
+Despite significant advances in techniques used for ECG inverse inference, uncertainty remains inadequately quantified across many aspects of the pipeline.
+Uncertainty quantification (UQ) aims to quantify the range of possible outcomes and their associated probabilities due to input variability [231].
+Recently, UQ has been widely employed in cardiac EP modeling, primarily for forward problems, to examine how variations in geometry and EP parameters affect the simulated ECG morphology [232, 233, 234].
+Instead, several studies have focused on how input variations inversely affect the reconstructed cardiac sources or estimated EP parameters, namely inverse UQ.
+These is generally achieved via polynomial chaos expansion, which allows for a compact and efficient representation of the uncertainty in model outputs [235, 236, 237].
+Furthermore, probabilistic approaches, which are often used for inverse UQ, naturally trace uncertainty from observed data back to the model estimation.
+ECG inverse problems are typically ill-posed, often lacking unique solutions, especially in complex scenarios such as activation reconstruction or model personalization incorporating CV and Purkinje networks.
+While uniqueness can sometimes be recovered in standard ECG imaging (the linear case), it is less likely in these more advanced models.
+However, clinicians typically require a single, reliable prediction for decision-making in patient care.
+Therefore, it is crucial to quantify and mitigate various uncertainties to enhance the robustness and clinical applicability of ECG inverse solutions.
+As illustrated in the ECG inverse inference pipeline presented in Fig.
+1, the primary sources of uncertainty include measurement errors, anatomical modeling inaccuracies, and assumptions in computational modeling.
+ECG or BSP data typically exhibit sparsity with measured values associated with a higher degree of uncertainty.
+For instance, Njeru et al.
+[238] quantified the uncertainty on the reconstructed EpiP due to the ECG measurement.
+Moreover, there is ambiguity in matching the BSPM measurement points with the computational mesh due to the non-synchronized BSPM and imaging acquisition [139].
+The cardiac position is also susceptible to various sources of movement artifacts, including respiration and cardiac contraction, introducing geometric errors into the ECGI solutions [239].
+The geometric error in the anatomical modeling (forward model) should also be considered, as it will in turn cause estimation errors in the ECG inverse inference [125, 130, 240].
+The anatomical modeling is normally started with image segmentation, so several studies directly performed UQ to investigate the effects of inaccurate segmentation on the reconstructed cardiac source [235, 241].
+Incorporating a prior model can mitigate uncertainties, and a combination of prior models may be preferred for estimating complex source structures [137].
+Also, one can disentangle the inter-subject anatomical variation to mitigate its effect on the ECG inverse problem [242].
+Besides the heart, the torso geometry accuracy can also subtly influence the inverse solution, as the EP measurement is generally performed on the body surface [243].
+There exist several studies performed UQ to investigate the effects of cardiac position inside the torso and torso conductivity on the inverse inference [244, 245, 237].
+To address uncertainties, advanced regularization techniques that encompass prior knowledge, noise reduction strategies, and advanced data assimilation methods show promise [54, 246].
+Model assumptions are necessary simplifications or idealizations of reality that are made to facilitate computational modeling.
+In the ECG inverse inference, assumptions are often made about the homogeneity of tissue properties, the behavior of electrical signals within the heart, and the relationship between measured ECG signals and underlying electrical activity [235, 247, 248].
+While model assumptions are essential for simplifying complex systems, they can also introduce uncertainties and errors into the modeling process.
+For example, assuming uniform tissue properties throughout the heart and torso may lead to inaccuracies in the predicted electrical activity, especially in regions with significant heterogeneity [243, 249].
+Similarly, assuming specific boundary conditions or physiological behaviors may not fully capture the complexity of real-world scenarios, leading to discrepancies between model predictions and observed data.
+ECG inverse inference also depends heavily on the assumptions on the initial distribution of heart electrical potential [231].
+Determining the initial potential distribution at the onset of the cardiac cycle is particularly challenging due to the difficulty in accurately pinpointing the anatomical site of the earliest excitation in the heart, often associated with the intricate Purkinje networks [117].
+Therefore, sophisticated modeling strategies are required to adequately account for the influence of these networks on ECG inverse inference [250, 194].
+Once uncertainties are quantified, strategies can be employed to mitigate their impact on model predictions.
+The process of balancing model assumptions and uncertainty is often iterative, requiring continuous refinement and validation of models against experimental or clinical data.
+This iterative refinement process helps researchers to build more robust and reliable models over time.
+5.2 Surrogate Model for Efficient ECG Simulation in Cardiac Digital Twin Development
+The functional twinning of cardiac EP often requires extensive simulations to achieve personalized parameter optimization.
 Traditional methods for simulating ECG signals involve complex biophysical models that account for intricate cardiac electrical and anatomical properties.
@@
 [252] presents a GPU-based implementation of the monodomain model, which is capable of speeding up simulations by a factor up to 498 when compared to a serial implementation of the model.
-They reported simulation times of 32 minutes per heartbeat.
+They reported simulation times of 32 min per heartbeat.
 Other studies have proposed alternative models that can yield similar simulation outputs at a fraction of the state-of-the-art monodomain models.
 For example, Wallman et al.
-[121] proposed a graph-based Eikonal model that can simulate the activation sequence in the heart in less than 1.6 seconds [103].
+[253] proposed a graph-based Eikonal model that can simulate the activation sequence in the heart in less than 1.6 s [1].
 Similarly, Neic et al.
-[213] presented a pseudo-monodomain model that can produce nearly identical simulations to a monodomain model (under sinus rhythm) but with a computational cost of 213 seconds per heartbeat [214].
+[254] introduced a reaction-Eikonal model, later implemented with a full pseudo-bidomain approach by Gillette et al.
+[2], which reported a computational time of 213 s per heartbeat.
+A simplified version using a lead field approach, without repolarization, can potentially be 10 times faster.
+These studies demonstrate a range of efficient ECG simulation tools, though reported times can vary significantly depending on model complexity (e.g., isotropic vs.
+anisotropic models), mesh resolution, and the hardware used in each approach [255, 2].
 Nevertheless, in the context of generating digital twins, the number of simulations required per patient can easily be in the range of hundreds of thousands [1, 194].
-Thus, even the fastest mechanistic models available present a significant computational burden when designing studies involving the generation of digital twins at scale [214, 138].
+Thus, even the fastest EP models available present a significant computational burden when applied at scale for digital twin generation [2, 194].
 The emergence of neural surrogate models addresses these challenges by offering computationally efficient alternatives that balance accuracy with practicality [119, 256, 257].
-These surrogate models leverage machine learning techniques to learn the intricate mapping between the physiological parameters (e.g., earliest activation sites) and expected behavior (e.g., cardiac activation sequences) [217].
+These surrogate models leverage machine learning techniques to learn the intricate mapping between the physiological parameters (e.g., EASs) and expected behavior (e.g., cardiac activation sequences) [257].
 For example, Bertrand et al.
-[217] was able to produce simulations of the activation sequence in 0.04 seconds, which was a 500-factor speed-up compared to a serial implementation of the graph-based Eikonal [121] (20 seconds per simulation).
+[257] was able to produce simulations of the activation sequence based on U-Net in 0.04 s, which was a 500-factor speed-up compared to a serial implementation of the graph-based Eikonal.
+Their surrogate EP model was generic and simplified (CV was fixed) and required an additional 5h 32 min of training time, but once trained, it can be deployed with high efficiency.
+Instead, Jiang et al.
+[256] developed a personalized neural surrogate model using meta-learning, reducing simulation time to 0.24 s and personalization time to 0.032 s, with an overall training time of about 8.5 min, significantly faster than the conventional 5 min AP simulation requiring 100 simulation calls.
 By circumventing the need for exhaustive simulations, surrogate models provide an avenue for rapid and on-demand generation of cardiac functional data (e.g., ECG signals).
@@
 Furthermore, a thorough understanding of the trade-off between computational efficiency and simulation accuracy is vital when integrating surrogate models into clinical applications.
-As surrogate models continue to evolve, they hold the potential to revolutionize ECG inverse inference by enabling researchers and practitioners to efficiently explore the parameter space and improve our understanding of complex cardiac phenomena.
+As surrogate models continue to evolve, they hold the potential to revolutionize CDTs by enabling researchers to efficiently explore the parameter space and improve our understanding of complex cardiac phenomena.
 5.3 Multimodal Representation Learning Towards Solving the ECG Inverse Problem
-The ECG inverse inference involves the fusion of information from multiple modalities, such as ECG signals, anatomical images (e.g., CT/ MRI scans), and clinical metadata.
-Multimodel representation learning offers a promising solution by allowing the integration of diverse data sources, leading to improved insights into the complex dynamics of cardiac electrical behavior.
-This can help address the issue of limited electrode coverage by inferring missing data points through learned relationships from other modalities.
-Furthermore, the inclusion of anatomical images can enhance the anatomical accuracy of the models, leading to improved source localization and more accurate inverse solutions.
-Indeed, deep neural networks have proven to be a highly effective strategy for ECG interpretation and analysis via spatiotemporal representation learning [47].
+The ECG inverse inference involves the fusion of information from multiple modalities, such as ECG signals, anatomical images (e.g., CT/ MRI scans), and EP parameters.
+Multimodal representation learning, a technique widely used in the computer vision field, holds significant potential for ECG inverse inference by enabling the seamless integration of diverse data sources [258].
+Deep neural networks have proven to be a highly effective strategy for ECG interpretation and analysis via spatiotemporal representation learning [259].
 However, most works only considered ECG data for analysis and only a few works employed joint analysis with other data sources [102, 147, 260, 188].
-Among these works, VAE has been widely employed to learn a latent representation of ECG and geometry [220], along with incorporating uncertainty modelling for inverse inference [116].
+Among these multimodal works, VAE has been widely employed to learn a latent representation of ECG and geometry [261, 188], along with incorporating uncertainty modeling for inverse inference [102].
 However, these studies merely concatenated the acquired features or utilized a shared representation, lacking a thorough explanation of how these features align and contribute to holistic data comprehension.
@@
 Balancing the complexity of the physical model with the capacity of the machine learning algorithm while also ensuring the interpretability of the resulting features will require careful consideration and innovative methodologies.
-6.3 The Role of Cardiac Digital Twins in Advancing in Silico Clinical Trials
-Cardiac digital twins are revolutionizing the landscape of in silico clinical trials, offering unprecedented opportunities to accelerate drug development [138].
-By creating virtual replicas of the individual patient’s heart, cardiac digital twins enable highly personalized simulations that capture the intricate dynamics of cardiac function and response to interventions [221].
-Traditional clinical trials often face limitations in recruiting participants from specific demographic or clinical groups, leading to challenges in generalizing findings to broader populations [222].
-With cardiac digital twins, researchers can create virtual patient cohorts that span a wide range of demographic, clinical, and pathological characteristics.
-This diversity allows for a more comprehensive exploration of drug effects and treatment outcomes across different patient subgroups, enhancing the reliability and generalizability of trial results.
-At the same time, in silico trials significantly reduce the time and cost associated with recruiting participants, conducting experiments, and analyzing results.
-This accelerated pace facilitates the rapid iteration and refinement of drug candidates, ultimately expediting the journey from discovery to market availability.
-Moreover, in silico trials enable researchers to explore a broader range of hypotheses and scenarios than would be feasible in traditional clinical settings.
-This flexibility allows for a more comprehensive exploration of potential drug effects, interactions, and mechanisms of action, leading to deeper insights and more informed decision-making.
-In the future, in silico clinical trials hold immense potential for revolutionizing drug development and healthcare delivery.
-The integration of patient-specific data and virtual patient models will enable personalized drug development approaches tailored to individual characteristics and needs.
-To fully realize the potential of in silico clinical trials, challenges such as model validity, regulatory compliance, and data access barriers must be addressed.
 6 Conclusion
-Cardiac digital twins are pivotal in advancing precision medicine and improving patient outcomes in cardiovascular care by simulating and predicting personalized responses to interventions.
-However, current models encounter a computational bottleneck in calibrating to an individual’s physiology through ECG inverse inference, particularly in 3D heart models.
-This review comprehensively outlines recent progress in ECG inverse inference, encompassing algorithms, public datasets, evaluation metrics, clinical applications, and future perspectives.
+CDTs are pivotal in advancing precision medicine and improving patient outcomes in cardiovascular care by simulating and predicting personalized responses to interventions.
+However, current models encounter a computational bottleneck in accurately calibrating CDTs to individual physiology through ECG inverse inference, particularly in estimating EP parameters within 3D heart models.
+This review comprehensively outlines recent progress in ECG inverse inference, encompassing algorithms, public datasets, evaluation, clinical applications, and future perspectives.
 While machine learning, especially deep learning, holds promise in tackling this complex and ill-posed problem, significant challenges persist, alongside promising opportunities for further advancement and innovation.
 In the future, as these techniques gain increasing popularity and generate massive cohorts of digital twins, they are poised to revolutionize cardiovascular care through in-silico trials and personalized interventions.
-With the proliferation of big data, envision a future where everyone possesses a digital twin, significantly reducing the reliance on animal models and enabling real-time monitoring in hospitals.
+With the growing availability of big data, we can envision a future where everyone possesses a digital twin, significantly reducing the reliance on animal models and enabling real-time monitoring and precision treatment in clinical settings.